### v3

# Tabular Data: Is Deep Learning all you need?

Abstract
Tabular data represent one of the most prevalent data formats in applied machine learning, largely because they accommodate a broad spectrum of real-world problems. Existing literature has studied many of the shortcomings of neural architectures on tabular data and has repeatedly confirmed the scalability and robustness of gradient-boosted decision trees across varied datasets. However, recent deep learning models have not been subjected to a comprehensive evaluation under conditions that allow for a fair comparison with existing classical approaches. This situation motivates an investigation into whether recent deep-learning paradigms outperform classical ML methods on tabular data. Our survey fills this gap by benchmarking seventeen state-of-the-art methods, spanning neural networks, classical ML and AutoML techniques. Our empirical results over 68 diverse datasets from a well-established benchmark indicate a paradigm shift, where Deep Learning methods outperform classical approaches.
1 Introduction
Tabular data has long been one of the most common and widely used data formats, with applications spanning various fields such as healthcare (Johnson et al., 2016; Ulmer et al., 2020), finance (Nureni & Adekola, 2022), and manufacturing (Chen et al., 2023), among others. Despite being a ubiquitous data modality, tabular data has only been marginally impacted by the deep learning revolution (Van Breugel & Van Der Schaar, 2024). A significant portion of the research community in tabular data continues to advocate for traditional machine learning methods, such as gradient-boosting decision trees (GBDTs) (Friedman, 2001; Chen & Guestrin, 2016; Prokhorenkova et al., 2018; Ke et al., 2017). Recent empirical studies suggest that GBDTs are still competitive for tabular data (Shwartz-Ziv & Armon, 2022; Grinsztajn et al., 2022; McElfresh et al., 2023). Nevertheless, an increasing segment of the community highlights the benefits of deep learning methods (Kadra et al., 2021; Gorishniy et al., 2021; Arik & Pfister, 2021; Somepalli et al., 2021; Kadra et al., 2024; Holzmüller et al., 2024).
The community remains divided on whether Deep Learning approaches are the undisputed state-of-the-art methods for tabular data (Shwartz-Ziv & Armon, 2022). To resolve this debate and determine the most effective methods for tabular data, multiple recent studies have focused on empirically comparing GBDTs with Deep Learning methods (Grinsztajn et al., 2022; Borisov et al., 2022; McElfresh et al., 2023). These studies suggest that tree-based models outperform deep learning models on tabular data even after tuning neural networks.
However, these recent empirical surveys only include non-meta-learned neural networks (Grinsztajn et al., 2022; Borisov et al., 2022) and do not incorporate the recent stream of methods that leverage foundation models and LLMs for tabular data (Zhu et al., 2023; Hollmann et al., 2023; Yan et al., 2024; Kim et al., 2024). Furthermore, the empirical setup of the recent empirical benchmarks is sub-optimal because no thorough hyperparameter optimization (HPO) techniques were applied to carefully tune the hyperparameters of neural networks.
In this empirical survey paper, we address a simple question: ”Is Deep Learning now state-of-the-art on tabular data, compared to GBDTs?”. Providing an unbiased and empirically justified answer to this question has a significant impact on the large community of practitioners. Therefore, we designed a large-scale experimental protocol using 68 diverse classification OpenML datasets and 17 recent baselines, including foundation models for tabular data. We classify models according to their underlying paradigm and provide a taxonomy tree in Figure 1. In our protocol, we use 10-fold cross-validation experiments for all the datasets and fairly tune the hyperparameters of all the baselines with an equally large HPO budget. In our study, we focus on the predictive quality of models rather than interpretability. However, recent works exist that propose interpretable counterparts for the top-performing deep learning methods in our study (Kadra et al., 2024; Mueller et al., 2024).
Figure 1: Taxonomy tree of algorithms applied to tabular classification (TC) models
Moreover, to fully unlock a model’s potential, contrary to prior work, after HPO we refit all models on the joined training and validation set. Hence, our study provides a fair investigation of post-hyperparameter optimization. We argue that this is a crucial oversight because training on the combined dataset can further improve a model’s predictive performance and change the ranking of the models, as indicated by our empirical results. Our findings highlight a paradigm shift, where Deep Learning methods achieve state-of-the-art results and manage to outperform classical approaches.
In summary, we provide the following main insights:
  • Meta-learned foundation models and simple feed-forward neural networks outperform GBDTs in all dataset regimes.
  • Feed-forward neural networks outperform related dataset-specific architectures, and non-fine-tuned foundation models outperform fine-tuned ones.
  • Refitting on the validation dataset after performing hyperparameter optimization significantly improves predictive quality and affects the overall model rankings.
  • As an overarching contribution, to facilitate future research, we open-source our code and we release a large benchmark that includes 17 baselines run on 68 datasets, repeated 10 times with different test outer folds, for up to 100 HPO trials, yielding a total number of 8 million evaluations or a total of 11.81 GPU years.
2 Related Work
Given the prevalence of tabular data in numerous areas, including healthcare, finance, psychology, and anomaly detection, as highlighted in various studies (Chandola et al., 2009; Johnson et al., 2016; Guo et al., 2017; Ulmer et al., 2020; Urban & Gates, 2021; Nureni & Adekola, 2022; Van Breugel & Van Der Schaar, 2024), there has been significant research dedicated to developing algorithms that effectively address the challenges inherent in these domains.
Classical Machine Learning.  Gradient Boosted Decision Trees (GBDTs) (Friedman, 2001), including popular implementations like XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), and CatBoost (Prokhorenkova et al., 2018), are widely favored by practitioners for their robust performance on tabular datasets and their short training times.
Deep Learning.  In terms of neural networks, prior work shows that meticulously searching for the optimal combination of regularization techniques in simple multilayer perceptrons (MLPs) called Regularization Cocktails (Kadra et al., 2021) can yield impressive results. Additionally, the models in (Kadra et al., 2021; Gorishniy et al., 2021) propose adaptations of the ResNet (He et al., 2016) architecture for tabular data, demonstrating the potential of deep learning approaches in handling tabular data.
Protocol Model families
Study Refitting Model-based HPO # Datasets GBDT AutoML ICL FT FNN TF # Baselines
Ye et al. (2025) × 300 × × 31
Rubachev et al. (2024) × 8 × × × 14
McElfresh et al. (2023) × × 176 × × 19
Borisov et al. (2022) × 5 × × × 20
Grinsztajn et al. (2022) × × 45 × × × 7
Shwartz-Ziv & Armon (2022) × 11 × × × × 5
Gorishniy et al. (2021) × 11 × × × 11
Ours 68 17
Table 1: Comparison with prior empirical survey works. In our study, we include 6 model families: Gradient Boosted Decision Trees (GBDT), AutoML, In-Context Learning (ICL), Fine-tuning (FT), Feed forward neural networks (FNN), and Transformer-based Models (TF).
Furthermore, recent research underscores that numerical embeddings (Gorishniy et al., 2022) for tabular data are underexplored. Incorporating these embeddings into neural network architectures, including MLPs and transformer-based models, can substantially enhance performance. Moreover, novel approaches such as RealMLP (Holzmüller et al., 2024) introduce various enhancements to the standard MLP architecture. These include using robust scaling at the preprocessing stage and experimenting with alternative numerical embedding strategies. Lastly, recent research (Gorishniy et al., 2025) achieves state-of-the-art performance by combining simple feed-forward neural networks with efficient ensembling techniques effectively mimicking gradient boosted decision trees.
Reflecting their success in various domains, transformers have also garnered attention in the tabular data domain. TabNet (Arik & Pfister, 2021), an innovative model in this area, employs attention mechanisms sequentially to prioritize the most significant features. SAINT (Somepalli et al., 2021) draws inspiration from the seminal transformer architecture (Vaswani et al., 2017). It addresses data challenges by applying attention both to rows and columns. They also offer a self-supervised pretraining phase, particularly beneficial when labels are scarce. The FT-Transformer (Gorishniy et al., 2021) stands out with its two-component structure: the Feature Tokenizer and the Transformer. The Feature Tokenizer is responsible for converting numerical and categorical features into embeddings. These embeddings are then fed into the Transformer, forming the basis for subsequent processing.
Recently, a new avenue of research has emerged, focusing on the use of foundation models for tabular data. XTab (Zhu et al., 2023) utilizes shared Transformer blocks, similar to those in FT-Transformer (Gorishniy et al., 2021), followed by fine-tuning dataset-specific encoders. Another notable work, TabPFN (Hollmann et al., 2023), employs in-context learning (ICL), by leveraging sequences of labeled examples provided in the input for predictions, thereby eliminating the need for additional parameter updates after training. The most recent version, TabPFNv2 (Hollmann et al., 2025), addresses the limitations of the first version, handling tables with up to 10K samples, and incorporating row- and column-wise attention, improving predictive performance. TabICL (Qu et al., 2025), similar to the TabPFN models, is pretrained on millions of synthetic datasets and can scale to tables with up to 500K samples. TP-BERTa (Yan et al., 2024), a pre-trained language model for tabular data prediction, uses relative magnitude tokenization to convert scalar numerical features into discrete tokens. The last layer of the model is then fine-tuned on a per-dataset basis. In contrast, CARTE (Kim et al., 2024) utilizes a graph representation of tabular data and a neural network capable of capturing the context within a table. The model is then fine-tuned on a per-dataset basis.
Empirical Studies.  Significant research has delved into understanding the contexts where neural networks (NNs) excel, and where they fall short (Shwartz-Ziv & Armon, 2022; Borisov et al., 2022; Grinsztajn et al., 2022; Rubachev et al., 2024; Ye et al., 2025). The recent study by McElfresh et al. (2023) is highly related to ours in terms of research focus. However, the authors use only random search for tuning the hyperparameters of neural networks, whereas we employ Tree-structured Parzen Estimator (TPE) (Bergstra et al., 2011) as employed by Gorishniy et al. (2021), which provides a more guided and efficient search strategy. Additionally, recent studies (McElfresh et al., 2023) are limited to evaluating a maximum of 30 hyperparameter configurations, in contrast to our more extensive exploration of up to 100 configurations. Furthermore, despite using the validation set for hyperparameter optimization (HPO), they do not retrain the model on the combined training and validation data using the best-found configuration before evaluating the model on the test set. Our paper differs from prior studies by applying a methodologically correct experimental protocol involving thorough HPO for neural networks. Lastly, Table 1 summarizes the model families evaluated in related empirical studies and highlights the differences in the evaluation protocol. To the best of our knowledge, we are the first to provide a thorough assessment of foundation models and AutoML to other learning paradigms.
3 Experimental Protocol
In our study, we focus on binary and multi-class classification problems on tabular data. The general learning task is described in Section 3.1. A detailed description of our evaluation protocol is provided in Section 3.2.
3.1 Learning with Tabular Data
A tabular dataset contains $N$ samples with $d$ features defining an $N\times d$ table. A sample $x_{i}\in\mathbb{R}^{d}$ is defined by its $d$ feature values. The features can be continuous numerical values or categorical, where for the latter, a common heuristic is to transform the values into numerical space. Given labels $y_{i}\in\mathcal{Y}$ being associated with the instances (rows) in the table, the task in our study is to solve a binary or multi-class classification problem. Hence, given a tabular dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, the aim is to learn a prediction model $f(\cdot)$ to minimize a classification loss function $\ell(\cdot,\cdot)$:
$\operatorname*{arg\,min}_{\theta}\sum_{(x_{i},y_{i})\in\mathcal{D}}\ell(y_{i},f(x_{i};\theta,\lambda)),$ (1)
where we use $f(x_{i};\theta,\lambda)$ for denoting the predicted label by a trained model parameterized by the model weights $\theta$ and hyperparameter configuration $\lambda$.
3.2 Experimental Setup
Datasets. In our study, we assess all the methods using OpenMLCC18 (Bischl et al., 2021), a well-established tabular benchmark in the community, which comprises 72 diverse datasets [^1] Due to memory issues encountered with several methods, we exclude four datasets from our analysis.. The datasets contain 5 to 3073 features and 500 to 100,000 instances, covering various binary and multi-class problems. The benchmark excludes artificial datasets, subsets or binarizations of larger datasets, and any dataset solvable by a single feature or a simple decision tree. For the full list of datasets used, we kindly refer the reader to Appendix E.
Preprocessing. We use a consistent preprocessing pipeline across all methods whenever possible. By default, we apply a quantile transformation using the scikit-learn library (Pedregosa et al., 2011), and categorical features are encoded with an ordinal encoder similar to prior work (Gorishniy et al., 2021). Methods for which we do not apply this preprocessing are those that inherently require a different approach, such as TP-BERTa and CARTE, or those implemented within libraries where modifying the preprocessing pipeline is not trivial. In these cases, we use the preprocessing strategies from the original works. Regarding batch size, we do not tune it in our experiments due to memory constraints. Instead, we determine batch size heuristically, similar to the setup proposed by Chen et al. (2024), based on the number of features in the dataset. While batch sizes may vary across datasets, they remain consistent across methods.
Evaluation Protocol. Our evaluation employs a nested cross-validation approach. Initially, we partition the data into 10 folds. Nine of these folds are then used for hyperparameter tuning. Each hyperparameter configuration is evaluated using 9-fold cross-validation. The results from the cross-validation are used to estimate the performance of the model under a specific hyperparameter configuration. For hyperparameter optimization, we utilize Optuna (Akiba et al., 2019), a well-known HPO library with the Tree-structured Parzen Estimator (TPE) (Bergstra et al., 2011) algorithm, the default Optuna HPO method. The optimization is constrained by a budget of either 100 trials or a maximum duration of 23 hours, similar to prior work (Kadra et al., 2021). Upon determining the optimal hyperparameters using Optuna, we train the model on the combined training and validation splits. All experiments are run on NVIDIA RTX2080Ti GPUs with 11 GB of memory. Our evaluation protocol dictates that for every algorithm, up to 68K different models will be evaluated, leading to a total of approximately 900K individual evaluations. As our study encompasses seventeen distinct methods, this methodology culminates in a substantial total of 8M evaluations, involving 900K unique models.
A detailed description of our evaluation protocol is provided in Appendix A.1. In our study, we adhere to the official hyperparameter search spaces from the respective papers for tuning every method. As a sole exception, the early stopping procedure is performed implicitly from the HPO procedure, where the number of training iterations is a hyperparameter similar to prior work (Kadra et al., 2021). We observed that this alternative form of early stopping yields better generalization. For a detailed description of the hyperparameter search spaces of all methods included in our analysis, we refer the reader to Appendix A.
Metrics. Lastly, we report the model’s performance as the average Area Under the Receiver Operating Characteristic (ROC-AUC) across 10 outer test folds. Given the prevalence of imbalanced datasets in the OpenMLCC18 benchmark, we employ ROC-AUC as our primary metric, since it offers a more reliable assessment of model performance.
Code: For reproducibility, our code is available at: https://github.com/machinelearningnuremberg/TabularStudy.
4 Baselines
In our experiments, we compare a range of methods categorized into three distinct groups: Classical Machine Learning Classifiers, Deep Learning Methods, and AutoML frameworks, as shown in Figure 1.
Classical Machine Learning Classifiers. First, we consider XGBoost (Chen & Guestrin, 2016), a well-established GBDT library that uses asymmetric trees. Moreover, we consider CatBoost (Prokhorenkova et al., 2018), a well-known library for GBDT that employs oblivious trees as weak learners and natively handles categorical features with various strategies. Finally, we also include LightGBM (Ke et al., 2017), a widely used GBDT framework that grows trees leaf-wise and supports efficient handling of large datasets.
Deep Learning Methods. In terms of classical deep learning methods, we include the ResNet implementation provided in the work by Gorishniy et al. (2021). Furthermore, we include three recent and competitive variants of the MLP architecture: i) an MLP architecture enhanced with numerical embeddings (Gorishniy et al., 2022) to which we refer as MLP, ii) RealMLP (Holzmüller et al., 2024), an MLP enhanced with several additions like robust scaling, numerical embeddings, etc, iii) TabM (Gorishniy et al., 2025), an efficient ensemble of MLP models.
In terms of transformer-based architectures, we consider TabNet (Arik & Pfister, 2021), an architecture that employs sequential attention to selectively utilize the most pertinent features at every decision step.
Next, we consider SAINT (Somepalli et al., 2021), a hybrid deep learning approach tailored for tabular data challenges. SAINT applies attention mechanisms across both rows and columns and integrates an advanced embedding technique. Lastly, we consider FT-Transformer (Gorishniy et al., 2021), an adaptation of the Transformer architecture for tabular data. It transforms categorical and numerical features into embeddings, which are then processed through a series of Transformer layers.
Foundation Models for Tabular Classification.  For in-context learning, we consider TabPFN (Hollmann et al., 2023), a meta-learned transformer architecture. Next, we consider TabPFNv2 (Hollmann et al., 2025), which alternates attention first across features, then across samples. Finally, we consider TabICL (Qu et al., 2025), pretrained on synthetic datasets similar to the TabPFN models, which can handle up to 500K samples. Among fine-tuned models, we include XTab (Zhu et al., 2023), a method that proposes a cross-table pretraining approach that can work across multiple tables with different column types and structures. Next, we consider TP-BERTa (Yan et al., 2024), a variant of the BERT language model that is adapted for tabular prediction. It introduces a relative magnitude tokenization to transform continuous numerical values into discrete high-dimensional tokens. Lastly, we include CARTE (Kim et al., 2024) in our experimental study. CARTE utilizes a graph representation of tabular data to process tables with differing structures.
Since all the fine-tuned models were pretrained on real-world datasets, we ensure that there are no datasets that overlap with the OpenMLCC18 benchmark.
AutoML Frameworks. Due to the large number of AutoML frameworks available in the community (Feurer et al., 2015; Erickson et al., 2020; LeDell & Poirier, 2020; Feurer et al., 2022), it was infeasible to include all of them in our experimental study. Therefore, we selected AutoGluon (Erickson et al., 2020), a framework that achieves the highest predictive performance in the recent AutoML Benchmark study (Gijsbers et al., 2024).
For all methods, we use their official implementations. We refer the readers to Appendix A for more details.
5 Experiments and Results
Figure 2: Left: Distribution of ranks for the Deep Learning ($12$ methods), Classical ML (33 methods) and AutoML (11 method) classifier families. Right: Distribution of ranks for the Foundation Models (55 methods), Dataset-Specific (77 methods) and AutoML (11 method) classifier families. The boxplots illustrate the rank spread, with medians represented by black lines, diamonds representing the means, and whiskers showing the range.
Research Question 1: Do DL models outperform gradient boosting methods in tabular data classification?
Figure 3: Win-rate dueling matrix comparing learning methods across shared datasets. Each cell (row $i$, column $j$) shows the fraction of common datasets on which method $i$ outperforms method $j$.
To address our research question, we initially compare the performance of Deep Learning methods and Classical ML methods jointly, by ranking the methods per-dataset and analyzing the rank distribution (the lower the rank, the better). The results provided in Figure 2 (Left) indicate that DL methods outperform the previous state-of-the-art GBDTs approaches. The best performing methods are TabICL with a median rank of 2, followed by AutoGluon and TabPFNv2, both with a median rank of 3. In terms of non-meta learned methods, TabM achieves a median rank of 5, followed by CatBoost, XGBoost, and LightGBM, with median ranks of 5.5, 7, and 7.5, respectively.
Next, we investigate how the models compare in a one-versus-one setting to eliminate the effect of related baselines. We present the results in Figure 3, where we observe that the one-versus-one results are consistent with the results where all the methods are considered jointly. In terms of meta-learned architectures, both TabPFNv2 and TabICL outperform tree-based architectures in the majority of the datasets. From the non-meta learned models, only TabM manages to outperform all variants of tree-based architectures.
Figure 4: Dataset landscape showing winning method families across different dataset sizes. Each point represents a dataset from the OpenMLCC18 benchmark, positioned by number of rows (x-axis) and features (y-axis) on log scales. Colors indicate which method family achieved the highest accuracy: Deep Learning methods (orange), Classical ML tree-based models (green), and ties (gray).
Lastly, we investigate whether there exists a certain region where deep learning methods are superior compared to the tree-based baselines, or where the opposite holds. For that purpose, we highlight the winning method family in Figure 4 for every dataset, over the number of examples and number of features of a dataset. The results show that Deep Learning methods dominate tree-based methods in datasets that have less than 5000 examples, by winning 31-3. In cases where a dataset has more than 5000 examples, tree-based methods become more competitive. However, they are still outperformed by deep learning methods 17-7.
For additional analyses, such as evaluations of predictive performance across different data regimes, examinations of the cost–efficiency trade-off, and investigations into the influence of meta-features on predictive performance, we refer readers to Appendices G, H, and I.
Research Question 2: Do meta-learned NNs outperform data-specific NNs in tabular data classification?
To answer the second research question, we analyze the distribution of ranks between the two subfamilies within the Deep Learning category: foundation models and dataset-specific neural networks. Figure 2 (Right) plot illustrates that in-context learning models are very competitive, with TabICL and TabPFNv2 having the best overall rank. Within the dataset-specific family, TabM demonstrates the best performance, attaining a median rank of 4 across all 68 datasets. The fine-tuned foundation models XTab, CARTE, and TP-BERTa obtain the worst performance compared to all remaining dataset-specific neural networks, except for TabNet. Interestingly to note is that, except for the in-context learning models, which are meta-learned architectures, the feed-forward neural networks TabM, RealMLP, and MLP outperform the attention-based architectures.
Next, we compare foundation models with dataset-specific NNs by generating critical difference diagrams. To generate the CD diagrams, we utilize the autorank package (Herbold, 2020), which performs a Friedman test followed by a Nemenyi post-hoc test at a significance level of $0.05$.
Figure 5: Critical difference (CD) diagram of the methods, where a horizontal bar indicates the absence of statistical significance. Left: CD diagram of Deep Learning vs. GBDTs, Right: CD diagram of dataset-specific vs. foundation models.
We present the results in Figure 5, where the black bars connecting methods indicate that there is no statistically significant difference in performance. Due to the limited number of datasets shared among the methods, TP-BERTa was excluded from this comparison. The left diagram of Figure 5 illustrates that TabPFNv2 and TabICL outperform all the other methods, demonstrating superior performance. TabM trails the top methods by a narrow margin, ranking third overall. CatBoost and XGBoost follow, and their performance differences relative to the top three are not statistically significant. We also compare the performance of models within the deep learning family. The right plot of Figure 5 shows a critical difference diagram indicating that TabICL and TabPFNv2 again attain the top average ranks. Except for TabM, both TabICL and TabPFNv2 are significantly better than all remaining methods.
A comprehensive presentation of the raw results for all methods, both after hyperparameter optimization (HPO) and with default hyperparameter configurations, is provided in Appendix D.
Figure 6: Left: Distribution of ranks for the Fine-tuning (3 models) and In-context learning (3 models) classifier families. The boxplots illustrate the rank spread, with medians represented by black lines and whiskers showing the range. Right: CD diagram of Fine-tuning models against In-context learning models. The horizontal bar indicates the absence of statistically significant differences.
Research Question 3: Which paradigm in transfer learning performs better: Do in-context models or fine-tuned models perform better?
To further investigate the family of foundation models, whether fine-tuning or in-context learning models yield better performance, we conducted an analysis similar to our previous research questions. We employ boxplots to display the distribution of ranks and use critical difference (CD) diagrams to evaluate the statistical significance of the results.
Figure 6 (Left) illustrates that TabICL, categorized under in-context learning methods, achieved a median rank of 1, and TabPFNv2 achieved a median rank of 2. Among the fine-tuning methods, XTab showed the best performance with a median rank of 3 but exhibited a larger interquartile range compared to the in-context learning methods, followed by CARTE and TP-BERTa.
To investigate whether the differences in performance are statistically significant, we present a CD diagram in Figure 6 (Right), from which TP-BERTa and TabPFN are excluded due to the limited number of common datasets among the methods. The CD diagram reveals that the in-context learning methods, TabICL and TabPFNv2, significantly outperform the fine-tuning methods XTab and CARTE.
Research Question 4: Does refitting after performing hyperparameter optimization have a significant impact on the predictive quality of the models, and does it impact the overall model ranking?
To investigate the impact of refitting, we select two top-performing distinct methods from the Deep Learning family, namely FT-Transformer as a transformer-based architecture, and TabM as an MLP-based architecture, while also selecting CatBoost and XGBoost from the tree-based family as the top-performing models.
Figure 7: Refitting impact on the predictive performance. Left: Investigating the rank distribution of the methods in isolation, with and without refitting. Right: Investigating the distribution of the ranks for the methods jointly, with and without refitting.
Initially, we compare the models in isolation to investigate how refitting affects the distribution of predictive performances across tasks. We present the results in Figure 7 (Left), where, as observed, all the methods that incorporate refitting feature a lower rank and outperform their non-refitting counterparts. Additionally, as we show in Appendix C, the difference in results is statistically significant in the majority of cases.
Moreover, we investigate whether refitting affects the ranking of the methods when considered jointly. The results in Figure 7 (Right) indicate that refitting does change the ranking of the methods, where, e.g., after refitting, XGBoost manages to outperform FT-Transformer and achieves a better median rank compared to the non-refitting counterpart. We continue by evaluating the impact of refitting via head-to-head, dataset-level comparisons, and applying statistical tests to assess the significance of the observed differences. We kindly refer the reader to Appendix C for details.
Research Question 5: What is the influence of hyperparameter optimization on a method’s predictive performance?
Figure 8: Rank improvement of methods with hyperparameter optimization compared to using the default configuration. Rankings are computed relative to other methods on each dataset (default rank – HPO rank; positive values indicate improvement). Box plots show the distribution across datasets, with points for individual datasets, horizontal bar for the median and diamonds for the mean.
To investigate the impact of hyperparameter optimization, we calculate the per-dataset rank of every method with and without performing hyperparameter optimization and compare the median rank improvement for every method compared to the other baselines. In general, the majority of methods improve in predictive performance when considered in isolation, as validated in Appendix B.13. However, the results in Figure 8 indicate that only TabM, RealMLP, SAINT, XGBoost, and XTab improve in terms of median rank compared to the other contenders, when hyperparameter optimization is performed.
In Appendix B, we provide a detailed analysis of hyperparameter importance, showing both the overall contribution of hyperparameters to model performance and the individual effect of each hyperparameter for every method.
6 Conclusion
Our comprehensive empirical study evaluates the quality of seventeen state-of-the-art tabular classification approaches across 68 diverse classification datasets from the OpenMLCC18 benchmark with a rigorous setup, employing cross-validation, model-based hyperparameter optimization, and refitting.
Our results indicate a paradigm shift, where deep learning methods outperform traditional baselines in all dataset regimes of the considered benchmark. Next to a fair comparison of model families, we provide an in-depth analysis of the importance of refitting, the influence of hyperparameter optimization on the models’ performance, the most important hyperparameters per method, and the cost-performance efficiency of various methods.
Our study contributes valuable insights into the current landscape of tabular data modeling and encourages further potential research directions with promising model families.
Acknowledgments
We acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under SFB 1597 (SmallData), grant number 499552394.
Appendix A Evaluation protocol and Configuration Spaces
A.1 Evaluation Protocol
1
Input : Dataset $D$, Number of outer folds $K=10$, Number of inner folds $J=9$, Number of hyperparameter optimization trials $T=100$, Search space $\Lambda$
Output : Overall performance $\bar{P}_{\text{outer}}$
2
3for $k\leftarrow 1$ to $K$ do
4   Split $D$ into training set $D_{\text{train}}^{k}$ and test set $D_{\text{test}}^{k}$;
5
6 for $t\leftarrow 1$ to $T$ do
7      Sample hyperparameter configuration $\theta_{t}$ from the search space $\Lambda$;
8
9    for $j\leftarrow 1$ to $J$ do
10         Split $D_{\text{train}}^{k}$ into inner training set $D_{\text{train}}^{k,j}$ and validation set $D_{\text{val}}^{k,j}$;
11         Train model $M(\lambda_{t})$ on $D_{\text{train}}^{k,j}$;
12         Evaluate $M(\lambda_{t})$ on $D_{\text{val}}^{k,j}$ to get performance $P^{k,j}(\lambda_{t})$;
13
14      end for
15
16     Compute mean performance $\bar{P}^{k}(\lambda_{t})=\frac{1}{J}\sum_{j=1}^{J}P^{k,j}(\lambda_{t})$;
17      Use $\bar{P}^{k}(\lambda_{t})$ as the objective value for $\lambda_{t}$;
18
19   end for
20
21  Select the best hyperparameter configuration $\lambda^{*}_{k}$ ;
22   Train final model $M(\lambda^{*}_{k})$ on $D_{\text{train}}^{k}$;
23   Evaluate $M(\lambda^{*}_{k})$ on $D_{\text{test}}^{k}$ to get outer performance $P_{\text{outer}}^{k}$;
24
25 end for
26
27Compute overall performance $\bar{P}_{\text{outer}}=\frac{1}{K}\sum_{k=1}^{K}P_{\text{outer}}^{k}$;
28 return $\bar{P}_{\text{outer}}$;
Algorithm 1 Nested Cross-Validation for Hyperparameter Optimization
Algorithm 1, shows the nested-cross validation with the outer folds (lines 1-1) and inner folds (lines 1-1). In each trial (lines 1-1), the mean performance across inner folds are calculated in line 1 which is used as the objective value for Optuna in line 1. After the maximal number of trials $T$ is reached or the time budget is exceeded, we select the best hyperparameter setting in line 1.
A.2 CatBoost
Table 2: Search space for CatBoost.
Parameter Type Range Log Scale
max_depth Integer [3, 10]
learning_rate Float $10^{-5}$, 1]     ✓
bagging_temperature Float [0, 1]
l2_leaf_reg Float [1, 10]     ✓
leaf_estimation_iterations Integer [1, 10]
iterations Integer [100, 2000]
The specific search space employed for CatBoost is detailed in Table 2. Our implementation heavily relies on the framework provided by the official implementation of the FT-Transformer, as found in the following repository [^2] https://github.com/yandex-research/rtdl-revisiting-models. We do this to ensure a consistent pipeline across all methods, that we compare. The CatBoost algorithm implementation, however, is the official one [^3] https://catboost.ai/.
For the default configuration of CatBoost, we do not modify any hyperparameter values. This approach allows the library to automatically apply its default settings, ensuring that our implementation is aligned with the most typical usage scenarios of the library.
A.3 XGBoost
Table 3: Search space for XGBoost.
Parameter Type Range Log Scale
max_depth Integer [3, 10]
min_child_weight Float $10^{-8}$, $10^{5}$]     ✓
subsample Float [0.5, 1]
learning_rate Float $10^{-5}$, 1]     ✓
colsample_bylevel Float [0.5, 1]
colsample_bytree Float [0.5, 1]
gamma Float $10^{-8}$, $10^{2}$]     ✓
reg_lambda Float $10^{-8}$, $10^{2}$]     ✓
reg_alpha Float $10^{-8}$, $10^{2}$]     ✓
n_estimators Integer [100, 2000]
We utilized the official XGBoost implementation [^4] https://xgboost.readthedocs.io/en/stable/. While the data preprocessing steps were consistent across all methods, a notable exception was made for XGBoost. For this method, we implemented one-hot encoding on categorical features, as XGBoost does not inherently process categorical values, in line with the implementation from the FT-Transformer repository.
The comprehensive search space for the XGBoost hyperparameters is detailed in Table 3. In the case of default hyperparameters, our approach mirrored the CatBoost implementation where we opted not to set any hyperparameters explicitly but instead, use the library defaults.
Furthermore, it is important to note that XGBoost lacks native support for the ROC-AUC metric in multiclass problems. To address this, we incorporated a custom ROC-AUC evaluation function. This function first applies a softmax to the predictions and then employs the ROC-AUC scoring functionality provided by scikit-learn, which can be found at the following link [^5] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.roc_auc_score.html.
A.4 LightGBM
Table 4: Search space for LightGBM.
Parameter Type Range Log Scale
feature_fraction Float [0.5, 1.0]
lambda_l2 Float {0.0, [0.1, 11.0]}     ✓
learning_rate Float [0.001, 1.0]     ✓
num_leaves Integer [4, 768]
min_sum_hessian_in_leaf float [0.0001, 100]     ✓
bagging_fractions Float [0.5, 1.0]
bagging_fractions Float [0.5, 1.0]
n_estimators Integer [100, 2000]
The hyperparameter search space for LightGBM is shown in Table 4. As with other methods, we adopt the preprocessing pipeline from the FT-Transformer repository.
For the default configuration, we retain all library-defined hyperparameters without modification.
A.5 FT-Transformer
Table 5: Search space for FT-Transformer.
Parameter Type Range Log Scale
n_layers Integer [1, 6]
d_token Integer [64, 512]
residual_dropout Float [0, 0.2]
attn_dropout Float [0, 0.5]
ffn_dropout Float [0, 0.5]
d_ffn_factor Float $\frac{2}{3}$, $\frac{8}{3}$]
lr Float $10^{-5}$, $10^{-3}$]     ✓
weight_decay Float $10^{-6}$, $10^{-3}$]     ✓
epochs Integer [10, 500]
In our investigation, we adopted the official implementation of the FT-Transformer (Gorishniy et al., 2021). Diverging from the approach from the original study, we implemented a uniform search space applicable to all datasets, rather than customizing the search space for each specific dataset. This approach ensures a consistent and comparable application across various datasets. The uniform search space we employed aligns with the structure proposed in Gorishniy et al. (2021). Specifically, we consolidated the search space by integrating the upper bounds defined in the original paper with the minimum bounds identified across different datasets.
Regarding the default hyperparameters, we adhered strictly to the specifications provided in Gorishniy et al. (2021).
A.6 SAINT
We utilize the official implementation of the method as detailed by the respective authors (Somepalli et al., 2021). The comprehensive search space employed for hyperparameter tuning is illustrated in Table 6.
Regarding the default hyperparameters, we adhere to the specifications provided by the authors in their original implementation.
Table 6: Search space for SAINT.
Parameter Type Range Log Scale
embedding_size Categorical {4, 8, 16, 32}
transformer_depth Integer [1, 4]
attention_dropout Float [0, 1.0]
ff_dropout Float [0, 1.0]
lr Float $10^{-5}$, $10^{-3}$]     ✓
weight_decay Float $10^{-6}$, $10^{-3}$]     ✓
epochs Integer [10, 500]
A.7 TabNet
Table 7: Search space for TabNet.
Parameter Type Range Log Scale
n_a Integer [8, 64]
n_d Integer [8, 64]
gamma Float [1.0, 2.0]
n_steps Integer [3, 10]
cat_emb_dim Integer [1, 3]
n_independent Integer [1, 5]
n_shared Integer [1, 5]
momentum Float [0.001, 0.4]     ✓
mask_type Categorical {entmax, sparsemax}
epochs Integer [10, 500]
For TabNet’s implementation, we utilized a well-maintained and publicly available version, accessible at the following link [^6] https://github.com/dreamquark-ai/tabnet. The hyperparameter tuning search space for TabNet, detailed in Table 7, was derived from  McElfresh et al. (2023).
Regarding the default hyperparameters, we followed the recommendations provided by the original authors.
A.8 ResNet
Table 8: Search space for ResNet.
Parameter Type Range Log Scale
layer_size Integer [64, 1024]
lr Float $10^{-5}$, $10^{-2}$]     ✓
weight_decay Float $10^{-6}$, $10^{-3}$]     ✓
residual_dropout Float [0, 0.5]
hidden_dropout Float [0, 0.5]
n_layers Integer [1, 8]
d_embedding Integer [64, 512]
d_hidden_factor Float [1.0, 4.0]
epochs Integer [10, 500]
We employed the ResNet implementation as described in prior work (Gorishniy et al., 2021). The entire range of hyperparameters explored for ResNet tuning is detailed in Table 8. Since the original study did not specify default hyperparameter values, we relied on the search space provided in a prior work (Kadra et al., 2021).
A.9 MLP-PLR
We employ the MLP implementation proposed by (Gorishniy et al., 2022). The search space used for hyperparameter optimization is detailed in Table 9. Default hyperparameters are adapted from (McElfresh et al., 2023), while the search space is based on the original work of (Gorishniy et al., 2022).
Table 9: Search space for MLP-PLR.
Parameter Type Range Log Scale
lr Float $10^{-5}$, $10^{-3}$]     ✓
weight_decay Float $10^{-6}$, $10^{-3}$]     ✓
dropout Float [0, 0.5]
n_layers Integer [1, 16]
d_embedding Integer [64, 512]
d_num_embedding Integer [1, 128]
d_first Integer [1, 1024]
d_middle Integer [1, 1024]
d_last Integer [1, 1024]
n Integer [1, 128]
sigma Float [0.01, 100]     ✓
epochs Integer [10, 500]
A.10 TabM
To run TabM in our experiments, we use the pytabkit implementation [^7] https://github.com/dholzmueller/pytabkit. The hyperparameter search space for TabM, presented in Table 10, is adapted from the original work (Gorishniy et al., 2025).
Table 10: Search space for TabM.
Parameter Type Range Log Scale
n_blocks Integer [1, 5]
d_block Integer [64, 1024]
dropout Float [0, 0.5]
hidden_dropout Float [0, 0.5]
lr Float $10^{-4}$, $5\times 10^{-3}$]     ✓
weight_decay Float $10^{-4}$, $10^{-1}$]     ✓
epochs Integer [10, 500]
A.11 RealMLP
For our RealMLP experiments, we use the official implementation in pytabkit7. Following the authors’ recommendations, we impute missing values using the mean of the training split before applying their preprocessing pipeline. We adopt the recommended default hyperparameters and search space, detailed in Table 11. Additionally, we extend the search space for initializating the standard deviation of the first embedding layer and tune the embedding dimensions, as suggested by the authors.
Table 11: Search space for RealMLP.
Parameter Type Range Log Scale
num_emb_type Categorical {None, PBLD, PL, PLR}
add_front_scale Categorical {True, False}
lr Float [2e-2, 3e-1]     ✓
p_drop Categorical {0.0, 0.15, 0.3}
act Categorical {relu, selu, mish}
hidden_sizes Categorical {[256, 256, 256], [64, 64, 64, 64, 64], [512]}
wd Categorical {0.0, 0.02}
plr_sigma Float [0.05, 1e1]     ✓
ls_eps Categorical {0.0, 0.1}
embedding_size Integer [1, 64]
n_epochs Integer [10, 500]
A.12 XTab
For XTab, we utilize the official implementation [^8] https://github.com/BingzhaoZhu/XTab. To ensure comparability with other methods, we decouple XTab from AutoGluon and apply the same preprocessing and training pipeline as used for the other models. The original work reports results for both light finetuning and heavy finetuning, so we introduce this as a categorical hyperparameter. If light_finetuning is set to True, the model is finetuned for only 3 epochs. Otherwise, we follow the same epoch range as for the other methods, i.e., $[10,500]$. Furthermore, we use the checkpoint after $2000$ iterations (iter_2k.ckpt), provided by the authors. Table 12 outlines the complete search space used for XTab during hyperparameter optimization.
Table 12: Search space for XTab.
Parameter Type Range Log Scale
lr Float $10^{-5},10^{-3}$]
weight_decay Float $10^{-6},10^{-3}$]
light_finetuning Categorical {True, False}
epochs Integer 33 (if light_finetuning=True) or $[10,500]$ (otherwise)
A.13 CARTE
For CARTE, we use the official implementation [^9] https://github.com/soda-inria/carte. Similar to XTab, since it is a pretrained model, we do not tune the architectural hyperparameters but keep them fixed and load the checkpoint provided by the authors. The search space used for CARTE during our hyperparameter optimization (HPO) process is shown in Table 13.
Table 13: Search space for CARTE.
Parameter Type Range Log Scale
lr Float $10^{-5},10^{-3}$]
weight_decay Float $10^{-6},10^{-3}$]
epochs Integer [$10,500$]
A.14 TP-BERTa
We use the official implementation for TP-BERTa [^10] https://github.com/jyansir/tp-berta. Similar to the other pretrained models, we only tune the learning rate, weight decay, and the number of finetuning epochs. The search space is shown in Table 14.
Table 14: Search space for TP-BERTa.
Parameter Type Range Log Scale
lr Float $10^{-5},10^{-3}$]
weight_decay Float $10^{-6},10^{-3}$]
epochs Integer [$10,500$]
A.15 TabPFN
For TabPFN and TabPFNv2, we utilized the official implementations from the authors [^11] https://github.com/automl/TabPFN. We followed the settings suggested by the authors and we did not preprocess the numerical features as TabPFN does that natively, we ordinally encoded the categorical features and we used an ensemble size of 32 for TabPFN to achieve peak performance as suggested by the authors. For TabPFNv2 we use the default settings and do not change anything.
A.16 TabICL
We use the official implementation of TabICL [^12] https://github.com/soda-inria/tabicl and follow the authors’ instructions without modification. In particular, we employ the default checkpoint tabicl-classifier-v1.1-0506.ckpt.
A.17 AutoGluon
For our experiments, we utilize the official implementation of AutoGluon [^13] https://auto.gluon.ai/stable/index.html. Specifically, we evaluate two configurations of AutoGluon: the HPO version and the recommended version.
  • For the HPO version, we use the default search spaces for the models included in AutoGluon’s ensemble.
  • For the recommended version, we set presets="best_quality" as per the official documentation and do not perform hyperparameter optimization.
Appendix B Hyperparameter analysis
In this section, we first examine the overall importance of hyperparameters for each method, as shown in Figure 9, which quantifies the contribution of each hyperparameter to model performance. The subsequent figures in this section illustrate the effect of individual hyperparameters on the performance metric. The x-axis represents the hyperparameters, while the y-axis denotes the ROC-AUC performance. We calculate hyperparameter importance using the fANOVA (Hutter et al., 2014) implementation in Optuna (Akiba et al., 2019). According to our analysis, the most important hyperparameter for CatBoost is the learning rate, while for XGBoost, it is the subsample ratio of the training instances. For XTab, the learning rate is also the most important hyperparameter, closely followed by the light_finetune hyperparameter, which is a categorical parameter taking values True or False. When light_finetune is True, we fine-tune XTab for only 3 epochs; when it is False, we use the same range of epochs as for the other methods (10 to 500). Similarly, for the MLP with PLR embeddings, the learning rate proves to be the most influential hyperparameter, whereas for RealMLP, the number of units in the hidden layers. For the remaining dataset-specific neural networks in the deep learning family, as well as for CARTE, the number of training epochs is the most important hyperparameter, indicating that training duration plays a critical role in their performance.
(a) CatBoost
(b) XGBoost
(c) ResNet
(d) FT-Transformer
(e) SAINT
(f) TabNet
(g) XTab
(h) CARTE
(i) TP-BERTa
(j) MLP
(k) RealMLP
(l) TabM
Figure 9: Hyperparameter importance for various methods
B.1 CatBoost
Figure 10: Effect of all the hyperparameters on model performance for CatBoost. The x-axis represents the hyperparameter values, while the y-axis shows the corresponding performance.
B.2 ResNet
Figure 11: Effect of all the hyperparameters on model performance for ResNet. The x-axis represents the hyperparameter values, while the y-axis shows the corresponding performance.
B.3 MLP-PLR
Figure 12: Effect of all the hyperparameters on model performance for MLP-PLR. The x-axis represents the hyperparameter values, while the y-axis shows the corresponding performance.
B.4 RealMLP
Figure 13: Effect of all the hyperparameters on model performance for RealMLP. The x-axis represents the hyperparameter values, while the y-axis shows the corresponding performance.
Since fANOVA does not support categorical hyperparameters, we exclude them from this analysis.
B.5 TabM
Figure 14: Effect of all the hyperparameters on model performance for TabM. The x-axis represents the hyperparameter values, while the y-axis shows the corresponding performance.
B.6 XGBoost
Figure 15: Effect of all the hyperparameters on model performance for XGBoost. The x-axis represents the hyperparameter values, while the y-axis shows the corresponding performance.
B.7 FT-Transformer
Figure 16: Effect of all the hyperparameters on model performance for FT-Transformer. The x-axis represents the hyperparameter values, while the y-axis shows the corresponding performance.
B.8 SAINT
Figure 17: Effect of all the hyperparameters on model performance for SAINT. The x-axis represents the hyperparameter values, while the y-axis shows the corresponding performance.
B.9 TabNet
Figure 18: Effect of all the hyperparameters on model performance for TabNet. The x-axis represents the hyperparameter values, while the y-axis shows the corresponding performance.
B.10 XTab
Figure 19: Effect of all the hyperparameters on model performance for XTab. The x-axis represents the hyperparameter values, while the y-axis shows the corresponding performance.
B.11 CARTE
Figure 20: Effect of all the hyperparameters on model performance for CARTE. The x-axis represents the hyperparameter values, while the y-axis shows the corresponding performance.
B.12 TP-BERTa
Figure 21: Effect of all the hyperparameters on model performance for TP-BERTa. The x-axis represents the hyperparameter values, while the y-axis shows the corresponding performance.
B.13 HPO influence on a per-model level
In our analysis of hyperparameter optimization (HPO) versus default configurations across various machine learning methods, we observed that HPO generally led to improved performance. The analysis is depicted in Figure 22. This improvement is reflected in the average rank reductions for most methods when HPO was applied. For example, XGBoost’s average rank improved significantly from 1.94 in its default configuration to 1.06 with HPO, and XTab showed a similar enhancement, moving from a rank of 1.96 down to 1.04.
These findings are visually represented in the accompanying plot, which illustrates the performance gains achieved through HPO. An exception to the general trend was observed with TP-BERTa, where the default configuration slightly outperformed the HPO version (average ranks of 1.47 and 1.53, respectively). This anomaly can be attributed to the computational demands of TP-BERTa. Due to its large model size, TP-BERTa was unable to complete the full 100 hyperparameter tuning trials within the allotted 23-hour time frame, often finishing only a few trials. Consequently, the HPO process may have converged to a suboptimal configuration that did not surpass the performance of the default settings.
Figure 22: Comparison of average rank performance between hyperparameter-optimized (HPO) models and default models. The blue dots represent the performance of the HPO models, while the red crosses denote the default models. Lower ranks indicate better performance.
Appendix C Ablating the Choice of Refitting
In this ablation, we explore whether refitting the model on the combined training and validation sets (after hyperparameter optimization) provides any measurable benefit. The standard procedure, as described in Section 3.2, uses a 10-fold nested cross-validation: we split the data into 10 folds, use 9 folds for inner cross-validation and HPO, then identify the best hyperparameter configuration and refit the model on all 9 folds before testing on the remaining fold.
We compare this approach to a no-refitting variant. Here, we still employ 10-fold cross-validation, but replace the inner cross-validation with a single 70/30 split of the 9 folds for training and validation. We train the model on the $70\%$ partition, perform HPO on the $30\%$ partition, and then save both the optimal hyperparameter configuration and the resulting trained model. Hence, when moving to the test fold, we simply load this trained model (with its fixed hyperparameters) instead of retraining on the entire 9-fold set. We repeat this for each of the 10 folds, ensuring the test set remains identical across both approaches.
Due to the computational resources required, we restrict this analysis to four methods: CatBoost, XGBoost, FT-Transformer, and TabM. The results of this ablation study are presented below, comparing performance with and without refitting.
Section 5, Research Question 4 (Figure 7) summarizes the main ablation results. Here, we extend that analysis by comparing the refitting and no-refitting variants head-to-head at the dataset level and conducting statistical tests to assess significance.
Figure 23 illustrates the performance difference between CatBoost with and without refitting across all datasets. The results clearly indicate that, with only a few exceptions, the refitted version consistently outperforms its non-refitted counterpart.
Figure 23: Performance difference between CatBoost with refitting and CatBoost without refitting across Datasets. Positive values indicate an improvement in ROC-AUC when refitting is applied, while negative values indicate a performance drop.
A similar pattern is observed in Figure 24 for XGBoost. Likewise, Figure 25 shows that refitting yields superior performance on most datasets for TabM too. In contrast, Figure 26 reveals that a greater number of datasets favor the non-refitted FT-Transformer. Nonetheless, the majority still benefit from refitting overall.
Figure 24: Performance difference between XGBoost with refitting and XGBoost without refitting across Datasets. Positive values indicate an improvement in ROC-AUC when refitting is applied, while negative values indicate a performance drop.
Figure 25: Performance difference between TabM with refitting and TabM without refitting across Datasets. Positive values indicate an improvement in ROC-AUC when refitting is applied, while negative values indicate a performance drop.
Figure 26: Performance difference between FT-Transformer with refitting and FT-Transformer without refitting Across Datasets. Positive values indicate an improvement in ROC-AUC when refitting is applied, while negative values indicate a performance drop.
Furthermore, we conducted a Wilcoxon signed-rank test to compare the performance of refitting versus no-refitting across multiple datasets for each method. The statistical results are summarized in Table 15.
Table 15: Statistical Comparison of Refit vs. No-Refit Methods
Method Pair #Datasets Avg. Diff Median Diff Wilcoxon Stat p-value CatBoost vs. CatBoost_noRefit 68 0.0079 0.0016 180.0000 1.298511e-09 XGBoost vs. XGBoost_noRefit 68 0.0144 0.0029 152.0000 4.413043e-10 TabM vs. TabM_noRefit 68 0.0056 0.0010 101.0000 3.395174e-10 FT vs. FT_noRefit 68 0.0035 0.0004 843.0000 9.356765e-02
For CatBoost, we observed an average performance improvement of $0.0079$ when refitting, with a median difference of $0.0016$. The Wilcoxon test yielded a test statistic of $180.0000$ and a highly significant p-value of $1.2985\cdot 10^{-9}$. This strongly suggests that refitting leads to a statistically significant and consistent improvement in CatBoost’s performance across datasets. Given the very low p-value $(p<0.001)$, we can confidently reject the null hypothesis that refitting has no effect.
For XGBoost, the average improvement with refitting was $0.0144$, with a median difference of $0.0029$. The Wilcoxon test statistic was $152.0000$, with a highly significant p-value of $4.4130\cdot 10^{-10}$. These results indicate that, similar to CatBoost, refitting yields a consistent and statistically significant improvement in performance for XGBoost.
In contrast, for FT-Transformer, the average improvement due to refitting was $0.0035$, with a median difference of $0.0004$. However, the Wilcoxon test yielded a test statistic of $843.0000$ and a p-value of $0.0936$, which is not statistically significant $(p>0.05)$. This suggests that while refitting improves FT-Transformer’s performance on average, the improvement is not consistent or significant across datasets.
For TabM, refitting led to an average performance improvement of $0.0056$ with a median difference of $0.0010$. The Wilcoxon test yielded a test statistic of $101.0000$ and a highly significant p-value of $3.3952\cdot 10^{-10}$. This provides strong evidence that refitting substantially and consistently improves TabM’s performance across datasets. Given the extremely low p-value $(p<0.001)$, we can confidently reject the null hypothesis, concluding that refitting has a significant positive impact on TabM.
Additionally, Table 16 presents the raw results of FT-Transformer, CatBoost, XGBoost and TabM, in comparison to their non-refitted counterparts.
Table 16: Average test ROC-AUC per dataset for FT, CatBoost, TabM and XGBoost using refitting vs. no refitting across CV folds.
Dataset CatBoost CatBoost_norefit XGBoost XGBoost_norefit FT FT_norefit TabM TabM_norefit adult 0.930747 0.924052 0.930482 0.916441 0.914869 0.915875 0.919662 0.917949 analcatdata_authorship 0.999662 0.999470 0.999816 0.999304 0.999985 0.999566 1.000000 1.000000 analcatdata_dmft 0.579136 0.547691 0.572150 0.575579 0.576947 0.579169 0.576017 0.568896 balance-scale 0.972625 0.962132 0.991268 0.889868 0.999735 0.995086 0.998912 0.998805 bank-marketing 0.938831 0.937464 0.938384 0.933394 0.938198 0.937470 0.941872 0.941008 banknote-authentication 0.999935 0.999979 0.999935 0.999849 1.000000 1.000000 1.000000 1.000000 Bioresponse 0.885502 0.872449 0.888615 0.853483 0.820159 N/A 0.876671 0.865593 blood-transfusion-service-center 0.754965 0.749848 0.750671 0.726849 0.745975 0.748119 0.748538 0.746875 breast-w 0.989162 0.992507 0.992112 0.991595 0.989503 0.989074 0.995845 0.995032 car 1.000000 0.998453 0.999902 0.999412 0.999751 0.999969 1.000000 1.000000 churn 0.922968 0.916146 0.914432 0.912343 0.914596 0.915300 0.929636 0.918852 climate-model-simulation-crashes 0.951480 0.944551 0.947000 0.947724 0.934671 0.933561 0.939969 0.940694 cmc 0.740149 0.735398 0.735649 0.721967 0.739402 0.736959 0.743797 0.738024 cnae-9 0.996316 0.994599 0.997454 0.983546 0.994497 0.994377 0.998100 0.997319 connect-4 0.921050 0.913372 0.931952 0.923175 0.901170 0.921978 0.941654 0.938588 credit-approval 0.934006 0.940661 0.934692 0.943379 0.935798 0.944236 0.934458 0.928719 credit-g 0.801762 0.773381 0.798571 0.794000 0.783048 0.777810 0.790905 0.770190 cylinder-bands 0.912070 0.867995 0.928116 0.866898 0.915494 0.826412 0.926477 0.886880 diabetes 0.837869 0.822365 0.835638 0.823473 0.831108 0.823379 0.829801 0.823632 dna 0.995028 0.994658 0.995278 0.994620 0.990937 0.989937 0.994505 0.994692 dresses-sales 0.595731 0.605008 0.622414 0.582184 0.620033 0.610016 0.642200 0.615764 electricity 0.980993 0.937421 0.987790 0.948766 0.963076 0.957884 0.968731 0.964049 eucalyptus 0.923334 0.916719 0.918055 0.902200 0.923933 0.911772 0.931897 0.920520 first-order-theorem-proving 0.831775 0.811589 0.834883 0.810288 0.796707 0.785106 0.818255 0.805200 GesturePhaseSegmentationProcessed 0.916674 0.779683 0.916761 0.774266 0.895166 0.799810 0.933828 0.908776 har 0.999941 0.999887 0.999960 0.999919 0.999685 0.999706 0.999966 0.999927 ilpd 0.744702 0.731536 0.748019 0.718251 0.751488 0.737753 0.744875 0.739189 Internet-Advertisements 0.979120 0.972513 0.982276 0.978762 0.974513 0.985391 0.985640 0.982167 isolet 0.999389 0.999282 0.999488 0.999225 0.998817 0.999282 0.999750 0.999628 jm1 0.756611 0.742362 0.759652 0.734592 0.709321 0.725904 0.751557 0.735722 jungle_chess_2pcs_raw_endgame_complete 0.976349 0.973983 0.974087 0.852209 0.999975 0.999861 0.999985 0.999945 kc1 0.825443 0.814042 0.832007 0.808270 0.783519 0.803310 0.813763 0.799767 kc2 0.846802 0.841593 0.843295 0.839860 0.832014 0.837281 0.833491 0.813933 kr-vs-kp 0.999392 0.999419 0.999796 0.999785 0.999777 0.999173 0.999652 0.999659 letter 0.999854 0.999802 0.999819 0.992828 0.999919 0.999886 0.999943 0.999906 madelon 0.937562 0.929178 0.932249 0.912814 0.747391 0.793476 0.809941 0.758538 mfeat-factors 0.998910 0.997917 0.999004 0.998767 0.999015 0.998560 0.999700 0.999550 mfeat-fourier 0.984714 0.984229 0.983375 0.981292 0.984511 0.982372 0.988497 0.987389 mfeat-karhunen 0.999264 0.998802 0.999211 0.998908 0.998682 0.997649 0.999521 0.999267 mfeat-morphological 0.965406 0.965867 0.963075 0.945833 0.970198 0.967869 0.969433 0.969514 mfeat-pixel 0.999422 0.999183 0.999378 0.998464 0.997451 0.998448 0.999478 0.999292 mfeat-zernike 0.977986 0.977831 0.974231 0.975436 0.983479 0.981858 0.984997 0.984036 MiceProtein 1.000000 0.999991 0.999923 0.999871 0.999973 1.000000 1.000000 0.999963 nomao 0.996439 0.995329 0.996676 0.995051 0.990908 0.992552 0.994828 0.994329 numerai28.6 0.529404 0.529350 0.529457 0.528295 0.530315 0.527963 0.529336 0.529882 optdigits 0.999844 0.999780 0.999855 0.999738 0.999616 0.999487 0.999939 0.999891 ozone-level-8hr 0.929094 0.923125 0.922663 0.918516 0.919484 0.919689 0.930601 0.926141 pc1 0.875471 0.850199 0.863061 0.853272 0.917591 0.840223 0.889312 0.862341 pc3 0.851122 0.833527 0.854543 0.833489 0.828743 0.835171 0.843468 0.839846 pc4 0.953309 0.945471 0.951037 0.903891 0.934944 0.944674 0.952956 0.950541 pendigits 0.999752 0.999728 0.999703 0.999777 0.999703 0.999668 0.999739 0.999756 PhishingWebsites 0.996482 0.995649 0.997425 0.996704 0.996760 0.996105 0.997636 0.996866 phoneme 0.968024 0.958699 0.967421 0.957712 0.965071 0.957862 0.971200 0.962861 qsar-biodeg 0.930649 0.928167 0.934875 0.917479 0.919584 0.914716 0.937730 0.929856 satimage 0.991978 0.990444 0.992114 0.990114 0.993516 0.992003 0.994291 0.992353 segment 0.996231 0.995441 0.996126 0.994624 0.994124 0.993598 0.994943 0.994696 semeion 0.998687 0.997784 0.998272 0.974216 0.995548 0.996208 0.998425 0.997714 sick 0.998331 0.997520 0.997950 0.995587 0.997937 0.997762 0.997317 0.996299 spambase 0.989935 0.988718 0.990726 0.988292 0.985969 0.983881 0.989244 0.988850 splice 0.995472 0.992511 0.995049 0.997548 0.992276 0.995195 0.995054 0.994709 steel-plates-fault 0.974350 0.968766 0.972743 0.954217 0.959182 0.962215 0.971043 0.966795 texture 0.999948 0.999946 0.999940 0.999834 0.999983 0.999973 0.999997 0.999997 tic-tac-toe 1.000000 0.999952 0.999710 0.999567 0.996152 0.996209 1.000000 1.000000 vehicle 0.943460 0.933394 0.942080 0.929008 0.963362 0.940233 0.965156 0.955308 vowel 0.999259 0.998833 0.999428 0.997587 0.999713 0.999198 0.999966 0.999854 wall-robot-navigation 0.999990 0.999910 0.999981 0.999586 0.999900 0.999870 0.999912 0.999873 wdbc 0.993813 0.991693 0.994467 0.993817 0.993967 0.986203 0.996573 0.994058 wilt 0.990950 0.991393 0.992192 0.991602 0.993047 0.992642 0.994857 0.995517
Appendix D Raw results tables
D.1 Results after hyperparameter optimization
Table 17 shows the raw results after HPO for CatBoost, LightGBM and XGBoost.
Table 17: Average test ROC-AUC per dataset for CatBoost, LightGBM and XGBoost after hyperparameter optimization across CV folds.
Dataset CatBoost LightGBM XGBoost adult 0.930747 0.931261 0.930482 analcatdata_authorship 0.999662 0.999986 0.999816 analcatdata_dmft 0.579136 0.573445 0.572150 balance-scale 0.972625 0.976799 0.991268 bank-marketing 0.938831 0.938470 0.938384 banknote-authentication 0.999935 0.999979 0.999935 Bioresponse 0.885502 0.886734 0.888615 blood-transfusion-service-center 0.754965 0.723635 0.750671 breast-w 0.989162 0.989162 0.992112 car 1.000000 0.999703 0.999902 churn 0.922968 0.913858 0.914432 climate-model-simulation-crashes 0.951480 0.950490 0.947000 cmc 0.740149 0.730807 0.735649 cnae-9 0.996316 0.984404 0.997454 connect-4 0.921050 0.932440 0.931952 credit-approval 0.934006 0.930848 0.934692 credit-g 0.801762 0.795774 0.798571 cylinder-bands 0.912070 0.929117 0.928116 diabetes 0.837869 0.827963 0.835638 dna 0.995028 0.994942 0.995278 dresses-sales 0.595731 0.617323 0.622414 electricity 0.980993 0.989807 0.987790 eucalyptus 0.923334 0.912027 0.918055 first-order-theorem-proving 0.831775 0.833949 0.834883 GesturePhaseSegmentationProcessed 0.916674 0.920034 0.916761 har 0.999941 0.999959 0.999960 ilpd 0.744702 0.711289 0.748019 Internet-Advertisements 0.979120 0.980094 0.982276 isolet 0.999389 0.999401 0.999488 jm1 0.756611 0.753206 0.759652 jungle_chess_2pcs_raw_endgame_complete 0.976349 0.977605 0.974087 kc1 0.825443 0.803456 0.832007 kc2 0.846802 0.843645 0.843295 kr-vs-kp 0.999392 0.999755 0.999796 letter 0.999854 0.999825 0.999819 madelon 0.937562 0.924095 0.932249 mfeat-factors 0.998910 0.999125 0.999004 mfeat-fourier 0.984714 0.983817 0.983375 mfeat-karhunen 0.999264 0.998958 0.999211 mfeat-morphological 0.965406 0.961250 0.963075 mfeat-pixel 0.999422 0.999131 0.999378 mfeat-zernike 0.977986 0.973627 0.974231 MiceProtein 1.000000 1.000000 0.999923 nomao 0.996439 0.996835 0.996676 numerai28.6 0.529404 0.529077 0.529457 optdigits 0.999844 0.999818 0.999855 ozone-level-8hr 0.929094 0.923040 0.922663 pc1 0.875471 0.874508 0.863061 pc3 0.851122 0.843094 0.854543 pc4 0.953309 0.946295 0.951037 pendigits 0.999752 0.999735 0.999703 PhishingWebsites 0.996482 0.997542 0.997425 phoneme 0.968024 0.965147 0.967421 qsar-biodeg 0.930649 0.931863 0.934875 satimage 0.991978 0.991309 0.992114 segment 0.996231 0.996233 0.996126 semeion 0.998687 0.997646 0.998272 sick 0.998331 0.998073 0.997950 spambase 0.989935 0.989995 0.990726 splice 0.995472 0.995103 0.995049 steel-plates-fault 0.974350 0.973788 0.972743 texture 0.999948 0.999890 0.999940 tic-tac-toe 1.000000 1.000000 0.999710 vehicle 0.943460 0.935611 0.942080 vowel 0.999259 0.998466 0.999428 wall-robot-navigation 0.999990 0.999968 0.999981 wdbc 0.993813 0.994729 0.994467 wilt 0.990950 0.985504 0.992192
Table 18 shows the raw results after HPO for dataset-specific neural networks.
Table 18: Average test ROC-AUC per dataset for dataset-specific neural networks after hyperparameter optimization across CV folds. Missing datasets are represented by ”-”.
Dataset FT-Transformer MLP RealMLP ResNet SAINT TabM TabNet adult 0.914869 0.928689 0.923327 0.913790 0.920246 0.919662 0.882450 analcatdata_authorship 0.999985 0.999770 1.000000 1.000000 0.999974 1.000000 0.999249 analcatdata_dmft 0.576947 0.574532 0.574396 0.584338 0.544695 0.576017 0.515962 balance-scale 0.999735 0.998659 1.000000 0.989061 0.999266 0.998912 0.979668 bank-marketing 0.938198 0.937054 0.937031 0.935740 0.936560 0.941872 0.887319 banknote-authentication 1.000000 1.000000 1.000000 1.000000 1.000000 1.000000 1.000000 Bioresponse 0.820159 0.825631 0.859065 0.850801 - 0.876671 - blood-transfusion-service-center 0.745975 0.770627 0.746350 0.738502 0.746726 0.748538 0.660675 breast-w 0.989503 0.992380 0.992882 0.995477 0.988470 0.995845 0.986694 car 0.999751 0.999992 1.000000 0.994154 1.000000 1.000000 1.000000 churn 0.914596 0.922938 0.913533 0.918713 0.915603 0.929636 0.891443 climate-model-simulation-crashes 0.934671 0.948857 0.962163 0.918990 0.925643 0.939969 0.868204 cmc 0.739402 0.744580 0.735472 0.737829 0.738490 0.743797 0.647121 cnae-9 0.994497 0.996716 0.997569 0.997106 - 0.998100 - connect-4 0.901170 0.927373 0.928258 0.933333 - 0.941654 - credit-approval 0.935798 0.938866 0.917352 0.933113 0.933493 0.934458 0.878500 credit-g 0.783048 0.788476 0.779381 0.783524 0.786402 0.790905 0.696905 cylinder-bands 0.915494 0.886405 0.910680 0.909989 0.923391 0.926477 0.837792 diabetes 0.831108 0.837342 0.837507 0.821798 0.827285 0.829801 0.756416 dna 0.990937 0.992220 0.994111 0.992543 0.992473 0.994505 0.991448 dresses-sales 0.620033 0.635468 0.537849 0.575205 0.624704 0.642200 0.555993 electricity 0.963076 0.969201 0.961467 0.960658 0.967012 0.968731 0.938656 eucalyptus 0.923933 0.921873 0.915693 0.916785 0.925970 0.931897 0.872365 first-order-theorem-proving 0.796707 0.798812 0.795637 0.784636 0.802392 0.818255 0.774094 GesturePhaseSegmentationProcessed 0.895166 0.911434 0.901441 0.914196 0.919006 0.933828 0.850596 har 0.999685 0.999783 0.999959 0.999921 - 0.999966 0.999515 ilpd 0.751488 0.671938 0.729412 0.747491 0.698718 0.744875 0.704840 Internet-Advertisements 0.974513 - 0.973810 0.974187 - 0.985640 - isolet 0.998817 0.998295 0.999635 0.999401 - 0.999750 0.998813 jm1 0.709321 0.715620 0.713988 0.720444 0.719464 0.751557 0.674043 jungle_chess_2pcs_raw_endgame_complete 0.999975 0.999965 0.999774 0.999956 0.999926 0.999985 0.991981 kc1 0.783519 0.805465 0.796117 0.806819 0.796918 0.813763 0.762807 kc2 0.832014 0.829426 0.845768 0.833248 0.834436 0.833491 0.713458 kr-vs-kp 0.999777 0.999686 0.999704 0.999369 0.999789 0.999652 0.998872 letter 0.999919 0.999894 0.999914 0.999926 0.999853 0.999943 0.999606 madelon 0.747391 0.883991 0.930302 0.605018 - 0.809941 0.630669 mfeat-factors 0.999015 0.998875 0.999625 0.999472 0.999385 0.999700 0.998125 mfeat-fourier 0.984511 0.984929 0.985483 0.981725 0.980508 0.988497 0.970539 mfeat-karhunen 0.998682 0.998849 0.999019 0.998448 0.999078 0.999521 0.996960 mfeat-morphological 0.970198 0.967719 0.969994 0.968651 0.967681 0.969433 0.955818 mfeat-pixel 0.997451 0.998674 0.999492 0.998690 0.999217 0.999478 0.998200 mfeat-zernike 0.983479 0.984610 0.982993 0.984488 0.981874 0.984997 0.968629 MiceProtein 0.999973 0.999973 0.999971 0.999973 1.000000 1.000000 0.999344 nomao 0.990908 0.986577 0.989803 0.993048 - 0.994828 - numerai28.6 0.530315 0.525920 0.529534 0.528012 0.525822 0.529336 - optdigits 0.999616 0.999794 0.999968 0.999927 0.999841 0.999939 0.998871 ozone-level-8hr 0.919484 0.927900 0.923252 0.925416 0.919315 0.930601 0.864067 pc1 0.917591 0.832532 0.844517 0.889458 0.870543 0.889312 0.804412 pc3 0.828743 0.842511 0.814590 0.829637 0.827322 0.843468 0.788151 pc4 0.934944 0.945813 0.939257 0.944447 0.934528 0.952956 0.920943 pendigits 0.999703 0.999705 0.999659 0.999638 0.999782 0.999739 0.999753 PhishingWebsites 0.996760 0.996991 0.997208 0.996975 0.996746 0.997636 0.996196 phoneme 0.965071 0.967617 0.966456 0.963591 0.960382 0.971200 0.956279 qsar-biodeg 0.919584 0.924951 0.929226 0.932220 0.930632 0.937730 0.902748 satimage 0.993516 0.992308 0.993034 0.991995 0.992630 0.994291 0.987482 segment 0.994124 0.995046 0.994075 0.993581 0.994831 0.994943 0.992317 semeion 0.995548 0.997350 0.998976 0.997689 0.997630 0.998425 0.994019 sick 0.997937 0.997048 0.998661 0.968841 0.998281 0.997317 0.981838 spambase 0.985969 0.988185 0.987799 0.987683 0.986263 0.989244 0.980804 splice 0.992276 0.994053 0.994420 0.993514 0.995073 0.995054 0.990441 steel-plates-fault 0.959182 0.964693 0.959639 0.949067 0.955379 0.971043 0.947456 texture 0.999983 0.999991 0.999999 0.999999 0.999976 0.999997 0.999763 tic-tac-toe 0.996152 1.000000 0.999711 0.999462 0.999725 1.000000 0.993030 vehicle 0.963362 0.961813 0.965844 0.967212 0.955127 0.965156 0.943787 vowel 0.999713 0.999638 0.999955 0.999813 0.999875 0.999966 0.999686 wall-robot-navigation 0.999900 0.999689 0.998720 0.999042 0.999844 0.999912 0.997585 wdbc 0.993967 0.996065 0.996038 0.995409 0.995546 0.996573 0.986656 wilt 0.993047 0.997690 0.993197 0.990726 0.993139 0.994857 0.991289
Table 19 shows the raw results after HPO for the meta-learned neural networks.
Table 19: Average test ROC-AUC per dataset for meta-learned neural networks after hyperparameter optimization across CV folds. Missing datasets are represented by ”-”.
Dataset CARTE TPBerta TabICL TabPFN TabPFNv2 XTab adult 0.902677 - 0.914430 - - - analcatdata_authorship 0.999181 - 1.000000 1.000000 1.000000 0.999991 analcatdata_dmft 0.586376 - 0.591336 0.586630 0.588236 0.556971 balance-scale 0.999413 - 0.997980 0.997656 0.995312 0.997420 bank-marketing 0.924664 - 0.940210 - - - banknote-authentication 1.000000 0.994512 1.000000 - 1.000000 1.000000 Bioresponse - - 0.885075 - - - blood-transfusion-service-center 0.739571 0.633041 0.743063 0.752586 0.754893 - breast-w 0.987912 0.986514 0.993152 0.994131 0.994132 0.989666 car 0.997126 - 0.999232 - 0.999963 - churn 0.923626 - 0.923984 - 0.923208 - climate-model-simulation-crashes 0.938531 - 0.932612 0.968010 0.958663 0.944367 cmc 0.738379 - 0.740035 - 0.746447 - cnae-9 0.990151 - 0.997840 - - - connect-4 - - 0.897904 - - - credit-approval 0.909279 0.901989 0.941488 0.932397 0.940813 0.939620 credit-g 0.769619 - 0.799048 0.768476 0.793429 - cylinder-bands 0.848539 0.820399 0.926679 0.886616 0.904451 0.881396 diabetes 0.823615 0.778356 0.835442 0.836120 0.844356 0.815847 dna 0.986120 - 0.994123 - 0.995658 0.992479 dresses-sales 0.589655 0.534893 0.605090 0.538916 0.608456 0.613136 electricity 0.909407 - 0.970809 - - 0.966899 eucalyptus 0.905245 - 0.934423 0.928493 0.933540 0.918317 first-order-theorem-proving 0.764092 - 0.834629 - 0.825502 0.798803 GesturePhaseSegmentationProcessed 0.798024 - 0.951408 - 0.936548 0.886960 har - - 0.999913 - - - ilpd 0.704712 0.586083 0.780714 0.757892 0.745193 0.726413 Internet-Advertisements - - 0.989308 - - - isolet - - 0.999570 - - - jm1 0.728512 - 0.784425 - - 0.727984 jungle_chess_2pcs_raw_endgame_complete 0.973383 - 0.975471 - - 0.999950 kc1 0.797680 - 0.849627 - 0.836795 0.803082 kc2 0.842828 - 0.834741 0.850065 0.837427 0.835476 kr-vs-kp 0.999685 0.855273 0.999792 - 0.999408 0.999616 letter 0.999440 - 0.999957 - - 0.999859 madelon 0.836760 - 0.711538 - - 0.845746 mfeat-factors 0.996064 - 0.999808 - 0.999650 0.998443 mfeat-fourier 0.976986 - 0.989372 - 0.991319 0.982539 mfeat-karhunen 0.994814 - 0.999850 - 0.999622 0.998582 mfeat-morphological 0.967325 - 0.968919 - 0.969308 0.967136 mfeat-pixel 0.996175 - 0.999664 - 0.999503 0.998642 mfeat-zernike 0.978119 - 0.992247 - 0.991483 0.980183 MiceProtein 0.999582 - 1.000000 - 1.000000 1.000000 nomao - - 0.996055 - - 0.992727 numerai28.6 0.514361 - 0.526838 - - 0.528062 optdigits 0.999112 - 0.999989 - 0.999897 0.999712 ozone-level-8hr 0.890063 - 0.936234 - 0.933398 0.915744 pc1 0.835444 - 0.912419 - 0.906419 0.855741 pc3 0.831574 0.625642 0.867955 - 0.854460 0.823532 pc4 0.937337 0.744304 0.959970 - 0.958887 0.938455 pendigits 0.999468 - 0.999852 - - 0.999751 PhishingWebsites 0.994582 - 0.998342 - - 0.996896 phoneme 0.948702 0.796404 0.977493 - 0.973546 0.961749 qsar-biodeg 0.921153 0.833852 0.943963 - 0.942895 0.926795 satimage 0.988038 - 0.993373 - 0.995122 0.992918 segment 0.993491 - 0.997462 - 0.997547 0.994697 semeion 0.993378 - 0.999033 - 0.998288 0.997064 sick 0.995762 - 0.997012 - 0.998008 0.998232 spambase 0.983228 - 0.992153 - 0.991218 0.986044 splice 0.987950 - 0.993905 - 0.995288 0.992444 steel-plates-fault 0.943636 - 0.978541 - 0.984454 0.957088 texture 0.999541 - 1.000000 - - 0.999962 tic-tac-toe 0.984361 0.993803 0.999182 0.996086 0.999663 1.000000 vehicle 0.941691 - 0.978088 0.970556 0.975896 0.955838 vowel 0.998092 - 1.000000 - - 0.999630 wall-robot-navigation 0.999505 - 0.999610 - 0.999936 0.999846 wdbc 0.990612 - 0.996697 0.996298 0.997761 0.994317 wilt 0.994858 0.880733 0.995557 - 0.996605 0.994261
Lastly, Table 20 shows the raw results of AutoGluon using HPO and AutoGluon with its recommended settings.
Table 20: Average test ROC-AUC per dataset for AutoGluon with HPO and AutoGluon with its recommended settings across CV folds.
Dataset AutoGluon AutoGluon (HPO) adult 0.931792 0.931658 analcatdata_authorship 1.000000 0.999887 analcatdata_dmft 0.577809 0.553672 balance-scale 0.997339 0.995057 bank-marketing 0.941273 0.940659 banknote-authentication 1.000000 0.999957 Bioresponse 0.888693 0.881238 blood-transfusion-service-center 0.741733 0.733305 breast-w 0.994394 0.993510 car 0.999861 0.999998 churn 0.927520 0.920213 climate-model-simulation-crashes 0.970051 0.926306 cmc 0.737077 0.536500 cnae-9 0.998524 0.997965 connect-4 0.934636 0.941976 credit-approval 0.940476 0.933497 credit-g 0.802381 0.773238 cylinder-bands 0.933320 0.903658 diabetes 0.833641 0.827171 dna 0.995385 0.994906 dresses-sales 0.615107 0.597537 electricity 0.987260 0.986609 eucalyptus 0.933782 0.925856 first-order-theorem-proving 0.835425 0.825561 GesturePhaseSegmentationProcessed 0.936667 0.917835 har 0.999958 0.999942 ilpd 0.765098 0.745564 Internet-Advertisements 0.985963 0.984740 isolet 0.999744 0.999696 jm1 0.770272 0.761065 jungle_chess_2pcs_raw_endgame_complete 0.999278 0.999444 kc1 0.835974 0.815660 kc2 0.834913 0.813625 kr-vs-kp 0.999405 0.999412 letter 0.999934 0.999933 madelon 0.932817 0.929882 mfeat-factors 0.999350 0.999111 mfeat-fourier 0.986058 0.986717 mfeat-karhunen 0.999575 0.998740 mfeat-morphological 0.977508 0.968908 mfeat-pixel 0.999403 0.999139 mfeat-zernike 0.995249 0.985279 MiceProtein 0.999929 0.999981 nomao 0.996892 0.996441 numerai28.6 0.530150 0.527692 optdigits 0.999925 0.999893 ozone-level-8hr 0.936029 0.930880 pc1 0.888177 0.860825 pc3 0.865766 0.845648 pc4 0.955384 0.950117 pendigits 0.999725 0.999642 PhishingWebsites 0.997572 0.997102 phoneme 0.973342 0.964555 qsar-biodeg 0.942988 0.932276 satimage 0.993557 0.993220 segment 0.996895 0.996421 semeion 0.998506 0.998210 sick 0.998367 0.997357 spambase 0.991092 0.989781 splice 0.995941 0.995249 steel-plates-fault 0.973843 0.972323 texture 0.999998 0.999995 tic-tac-toe 1.000000 0.996585 vehicle 0.969797 0.965886 vowel 0.999910 0.999618 wall-robot-navigation 0.999993 0.999984 wdbc 0.995799 0.992456 wilt 0.995652 0.994495
D.2 Results using default hyperparameter configurations
Table 21 shows the raw results for CatBoost and XGBoost using the default hyperparameter configurations.
Table 21: Average test ROC-AUC per dataset for CatBoost, LightGBM and XGBoost using the default hyperparamater configurations across CV folds.
Dataset CatBoost LightGBM XGBoost adult 0.930571 0.929995 0.929316 analcatdata_authorship 0.999710 0.999970 0.999518 analcatdata_dmft 0.549171 0.538902 0.531850 balance-scale 0.952530 0.920593 0.926923 bank-marketing 0.938725 0.937425 0.934864 banknote-authentication 0.999957 0.999613 0.999914 Bioresponse 0.879217 0.880857 0.880176 blood-transfusion-service-center 0.729842 0.706352 0.712258 breast-w 0.991254 0.990699 0.990430 car 0.999509 0.999672 0.998790 churn 0.924606 0.917109 0.913882 climate-model-simulation-crashes 0.962296 0.949276 0.955828 cmc 0.709590 0.695848 0.684939 cnae-9 0.996007 0.983430 0.994232 connect-4 0.893587 0.886247 0.899588 credit-approval 0.937424 0.925672 0.930615 credit-g 0.800667 0.787000 0.788381 cylinder-bands 0.885160 0.907731 0.912564 diabetes 0.835137 0.798912 0.797009 dna 0.994641 0.994798 0.994699 dresses-sales 0.598768 0.565517 0.570699 electricity 0.958153 0.954700 0.971787 eucalyptus 0.921691 0.903510 0.902805 first-order-theorem-proving 0.826532 0.828733 0.826895 GesturePhaseSegmentationProcessed 0.898407 0.889753 0.892459 har 0.999899 0.999938 0.999905 ilpd 0.741153 0.745050 0.722052 Internet-Advertisements 0.979992 0.978933 0.976972 isolet 0.999407 0.999095 0.998854 jm1 0.748060 0.749102 0.729353 jungle_chess_2pcs_raw_endgame_complete 0.972286 0.971617 0.974856 kc1 0.823661 0.791316 0.791182 kc2 0.821163 0.787678 0.771390 kr-vs-kp 0.999521 0.999727 0.999720 letter 0.999740 0.999724 0.999648 madelon 0.928172 0.902450 0.890107 mfeat-factors 0.999031 0.998867 0.998356 mfeat-fourier 0.984181 0.981478 0.982669 mfeat-karhunen 0.999128 0.998500 0.997700 mfeat-morphological 0.962489 0.955839 0.958908 mfeat-pixel 0.999289 0.998861 0.998703 mfeat-zernike 0.972961 0.965408 0.966633 MiceProtein 0.999983 0.999944 0.999680 nomao 0.995620 0.995122 0.995690 numerai28.6 0.518341 0.521861 0.511976 optdigits 0.999808 0.999690 0.999586 ozone-level-8hr 0.925485 0.916990 0.911594 pc1 0.891257 0.874492 0.857895 pc3 0.850219 0.819425 0.816916 pc4 0.953689 0.949308 0.942808 pendigits 0.999764 0.999772 0.999760 PhishingWebsites 0.995801 0.996155 0.996764 phoneme 0.955202 0.956695 0.957311 qsar-biodeg 0.934769 0.933991 0.926970 satimage 0.991815 0.990983 0.990907 segment 0.996012 0.996016 0.995267 semeion 0.998163 0.996984 0.996029 sick 0.998355 0.998355 0.996943 spambase 0.989066 0.990138 0.988888 splice 0.995198 0.994463 0.994788 steel-plates-fault 0.972233 0.973473 0.970148 texture 0.999908 0.999856 0.999795 tic-tac-toe 1.000000 0.998990 0.999181 vehicle 0.942832 0.936072 0.935079 vowel 0.999237 0.998215 0.996947 wall-robot-navigation 0.999989 0.999955 0.999934 wdbc 0.994217 0.992350 0.994471 wilt 0.991488 0.987326 0.988659
Table 22 shows the raw results for dataset-specific neural networks using the default hyperparameter configurations.
Table 22: Average test ROC-AUC per dataset for dataset-specific neural networks using default hyperparameter configurations across CV folds. Missing datasets are represented by ”-”.
Dataset FT-Transformer MLP RealMLP ResNet SAINT TabM TabNet adult 0.893029 0.897504 0.909085 0.905838 0.870099 0.908670 0.912781 analcatdata_authorship 0.999392 0.999934 0.999952 1.000000 0.999983 1.000000 0.993186 analcatdata_dmft 0.553755 0.554240 0.575007 0.553675 0.526597 0.539602 0.534271 balance-scale 0.988863 0.995111 0.980107 0.992229 0.991970 0.998859 0.972816 bank-marketing 0.907667 0.904699 0.814657 0.926617 0.892316 0.931979 0.927765 banknote-authentication 1.000000 1.000000 1.000000 1.000000 1.000000 1.000000 1.000000 Bioresponse 0.804580 0.560952 0.824996 0.843462 - 0.872512 0.812061 blood-transfusion-service-center 0.713181 0.762080 0.746119 0.742088 0.723673 0.727600 0.728919 breast-w 0.988615 0.994222 0.993411 0.991140 0.992220 0.995038 0.984383 car 0.999758 0.999678 1.000000 0.998600 0.999828 1.000000 0.931659 churn 0.915966 0.903166 0.917916 0.914732 0.910996 0.927037 0.905642 climate-model-simulation-crashes 0.840724 0.935571 0.947857 0.904025 0.937306 0.946051 0.825571 cmc 0.686016 0.710134 0.700557 0.687757 0.642394 0.692772 0.689043 cnae-9 0.994801 0.500463 0.992911 0.996595 - 0.997415 0.912423 connect-4 0.922969 0.915051 0.909829 0.926041 0.756318 0.938629 0.856762 credit-approval 0.915482 0.931215 0.914193 0.916769 0.908623 0.920201 0.875614 credit-g 0.731714 0.726875 0.758571 0.735071 0.744000 0.782714 0.632571 cylinder-bands 0.908565 0.874205 0.904906 0.891759 0.909314 0.924863 0.710240 diabetes 0.755846 0.829853 0.822211 0.789923 0.737127 0.789142 0.785077 dna 0.988362 0.990128 0.988320 0.992218 0.520670 0.993741 0.962713 dresses-sales 0.571921 0.536782 0.525944 0.536617 0.568144 0.542365 0.560591 electricity 0.963347 0.950665 0.950555 0.930924 0.960991 0.959880 0.911419 eucalyptus 0.917340 0.922173 0.903412 0.897582 0.904708 0.925170 0.877684 first-order-theorem-proving 0.796282 0.782461 0.781809 0.793079 0.772449 0.814144 0.743350 GesturePhaseSegmentationProcessed 0.827939 0.819054 0.890444 0.853272 0.893255 0.874995 0.781506 har 0.999876 0.999848 0.999630 0.999859 - 0.999937 0.999147 ilpd 0.724591 0.748217 0.727899 0.758030 0.713191 0.759658 0.715948 Internet-Advertisements 0.973465 0.982883 0.961953 0.967077 - 0.981634 0.892480 isolet 0.999463 0.847095 0.999135 0.999307 - 0.999671 0.997706 jm1 0.723314 0.726646 0.721977 0.734238 0.652524 0.738839 0.722615 jungle_chess_2pcs_raw_endgame_complete 0.998738 0.998486 0.996257 0.977410 0.999876 0.998544 0.974173 kc1 0.804719 0.801565 0.806604 0.795200 0.742990 0.820436 0.792858 kc2 0.805644 0.840419 0.829826 0.771497 0.742400 0.818842 0.806986 kr-vs-kp 0.999792 0.999765 0.998737 0.999476 0.723052 0.999796 0.987183 letter 0.999825 0.999640 0.999820 0.999864 0.999784 0.999918 0.997271 madelon 0.770769 0.500000 0.915592 0.600713 - 0.758018 0.559015 mfeat-factors 0.998765 0.998668 0.999075 0.998892 0.499849 0.999589 0.993717 mfeat-fourier 0.977475 0.978653 0.974028 0.980419 0.971772 0.986450 0.961111 mfeat-karhunen 0.997503 0.998582 0.999439 0.998097 0.998387 0.998958 0.982592 mfeat-morphological 0.967733 0.965494 0.968706 0.969308 0.967478 0.967408 0.963611 mfeat-pixel 0.997658 0.946632 0.999500 0.998676 0.553414 0.999317 0.992500 mfeat-zernike 0.978039 0.980681 0.965872 0.980858 0.969257 0.978049 0.966992 MiceProtein 1.000000 0.999963 1.000000 0.999963 1.000000 0.999991 0.987043 nomao 0.992049 0.991436 0.983015 0.992530 0.499521 0.994784 0.991441 numerai28.6 0.507813 0.513601 0.522412 0.517071 0.507780 0.523854 0.522797 optdigits 0.999631 0.999454 0.999927 0.999837 0.999057 0.999927 0.998476 ozone-level-8hr 0.893747 0.906572 0.822254 0.826296 0.881560 0.921495 0.869228 pc1 0.852119 0.853077 0.828996 0.820008 0.866325 0.891588 0.863233 pc3 0.810311 0.784672 0.768438 0.771759 0.804479 0.832408 0.809443 pc4 0.944764 0.940799 0.906347 0.936765 0.931286 0.949946 0.900752 pendigits 0.999740 0.999687 0.999850 0.999691 0.999785 0.999748 0.999088 PhishingWebsites 0.996882 0.996479 0.994417 0.997134 0.996805 0.997615 0.993856 phoneme 0.956543 0.948168 0.952913 0.938565 0.956949 0.963998 0.933545 qsar-biodeg 0.916158 0.924529 0.911174 0.916804 0.918103 0.934609 0.893489 satimage 0.992141 0.990975 0.986944 0.990613 0.985874 0.993552 0.986280 segment 0.994709 0.993795 0.994189 0.993821 0.993989 0.994829 0.992101 semeion 0.995507 0.968306 0.998289 0.996745 0.576269 0.998174 0.957550 sick 0.997877 0.989590 0.976784 0.969015 0.991121 0.995715 0.929353 spambase 0.983325 0.983168 0.978382 0.985056 0.981111 0.988298 0.978240 splice 0.989898 0.990919 0.991318 0.990917 0.991932 0.995071 0.972882 steel-plates-fault 0.959626 0.963250 0.955133 0.959356 0.948021 0.966012 0.916561 texture 0.999976 0.999956 0.999992 0.999999 0.996944 0.999997 0.999441 tic-tac-toe 0.998605 0.999145 0.997548 0.999375 0.996921 0.999904 0.899715 vehicle 0.956404 0.944588 0.961117 0.963268 0.944376 0.962549 0.923325 vowel 0.999618 0.997520 0.999641 0.999966 0.999888 0.999854 0.986644 wall-robot-navigation 0.999757 0.999245 0.998582 0.998972 0.999104 0.999520 0.997972 wdbc 0.994847 0.994219 0.998021 0.997080 0.997234 0.997765 0.985323 wilt 0.994235 0.994105 0.993080 0.994057 0.988766 0.994455 0.991840
Table 23 shows the raw results for the meta-learned neural networks using the default hyperparameter configurations.
Table 23: Average test ROC-AUC per dataset for meta-learned neural networks using default hyperparameter configurations across CV folds. Missing datasets are represented by ”-”.
Dataset CARTE TPBerta TabPFN XTab adult 0.897259 - - - analcatdata_authorship 0.998103 - 1.000000 0.997620 analcatdata_dmft 0.572113 - 0.586630 0.550627 balance-scale 0.998116 - 0.997656 0.895083 bank-marketing 0.907972 - - - banknote-authentication 1.000000 0.997535 - 0.996615 blood-transfusion-service-center 0.705189 0.659754 0.752586 - breast-w 0.984775 0.967673 0.994131 0.988527 car 0.992862 - - - churn 0.920360 - - - climate-model-simulation-crashes 0.938031 - 0.968010 0.568735 cmc 0.730370 - - - cnae-9 0.986921 - - - connect-4 0.500681 - - - credit-approval 0.906552 0.891294 0.932397 0.922447 credit-g 0.700952 - 0.768476 - cylinder-bands 0.810318 0.814857 0.886616 0.778646 diabetes 0.755348 0.768974 0.836120 0.822370 dna 0.981979 - - 0.992857 dresses-sales 0.591297 0.565189 0.538916 0.585057 electricity 0.874950 - - 0.900765 eucalyptus 0.907418 - 0.928493 0.814121 first-order-theorem-proving 0.735870 - - 0.721997 GesturePhaseSegmentationProcessed 0.771707 - - 0.737155 har - - - 0.999241 ilpd 0.729851 0.672431 0.757892 0.724427 isolet 0.995113 - - 0.998455 jm1 0.704730 - - 0.721445 jungle_chess_2pcs_raw_endgame_complete 0.918894 - - 0.965961 kc1 0.805108 - - 0.793122 kc2 0.826925 - 0.850065 0.835398 kr-vs-kp 0.958715 0.999107 - 0.995940 letter 0.998939 - - 0.989493 madelon 0.789929 - - 0.689657 mfeat-factors 0.794171 - - 0.997867 mfeat-fourier 0.969911 - - 0.956494 mfeat-karhunen 0.978967 - - 0.990728 mfeat-morphological 0.961442 - - 0.948069 mfeat-pixel 0.759099 - - 0.997478 mfeat-zernike 0.964453 - - 0.965907 MiceProtein 0.986177 - - 0.972404 nomao 0.981817 - - 0.991110 numerai28.6 0.521094 - - 0.527797 optdigits 0.998452 - - 0.999031 ozone-level-8hr 0.861468 - - 0.915294 pc1 0.791339 - - 0.729942 pc3 0.784448 0.683751 - 0.816464 pc4 0.907759 0.699487 - 0.888728 pendigits 0.999522 - - 0.999222 PhishingWebsites 0.991886 - - 0.987949 phoneme 0.932082 0.798855 - 0.911417 qsar-biodeg 0.914703 0.817997 - 0.919134 satimage 0.982299 - - 0.982955 segment 0.992163 - - 0.974072 semeion 0.983218 - - 0.989977 sick 0.991907 - - 0.950283 spambase 0.748573 - - 0.982966 splice 0.701980 - - 0.991116 steel-plates-fault 0.925718 - - 0.848468 texture 0.993709 - - 0.999521 tic-tac-toe 0.861176 0.958328 0.996086 0.744202 vehicle 0.929483 - 0.970556 0.893891 vowel 0.995589 - - 0.812581 wall-robot-navigation 0.998981 - - 0.986489 wdbc 0.993948 - 0.996298 0.984744 wilt 0.994112 0.960758 - 0.979966
Lastly, Table 24 shows the raw results of AutoGluon using the default settings.
Table 24: Average test ROC-AUC per dataset for AutoGluon using default configurations across CV folds.
Dataset AutoGluon adult 0.931179 analcatdata_authorship 0.999782 analcatdata_dmft 0.584732 balance-scale 0.594936 bank-marketing 0.939889 banknote-authentication 0.999957 Bioresponse 0.884276 blood-transfusion-service-center 0.741962 breast-w 0.992231 car 0.999593 churn 0.922201 climate-model-simulation-crashes 0.957745 cmc 0.691344 cnae-9 0.997878 connect-4 0.936000 credit-approval 0.935450 credit-g 0.783286 cylinder-bands 0.900459 diabetes 0.821997 dna 0.994904 dresses-sales 0.586043 electricity 0.987262 eucalyptus 0.754274 first-order-theorem-proving 0.830805 GesturePhaseSegmentationProcessed 0.920355 har 0.999938 ilpd 0.737184 Internet-Advertisements 0.984077 isolet 0.999636 jm1 0.764863 jungle_chess_2pcs_raw_endgame_complete 0.992186 kc1 0.821507 kc2 0.812567 kr-vs-kp 0.999619 letter 0.999901 madelon 0.925627 mfeat-factors 0.999142 mfeat-fourier 0.984642 mfeat-karhunen 0.998693 mfeat-morphological 0.969200 mfeat-pixel 0.998731 mfeat-zernike 0.982779 MiceProtein 0.899990 nomao 0.996397 numerai28.6 0.527789 optdigits 0.999670 ozone-level-8hr 0.927357 pc1 0.876676 pc3 0.849770 pc4 0.952137 pendigits 0.999684 PhishingWebsites 0.997256 phoneme 0.966521 qsar-biodeg 0.931279 satimage 0.992096 segment 0.996333 semeion 0.998341 sick 0.997864 spambase 0.989571 splice 0.995584 steel-plates-fault 0.971070 texture 0.999996 tic-tac-toe 0.999951 vehicle 0.958256 vowel 0.999641 wall-robot-navigation 0.898793 wdbc 0.992978 wilt 0.994524
Appendix E Datasets
In Table 25 we show a summary of all the OpenMLCC18 datasets used in this study.
Table 25: Summary of OpenML-CC18 Datasets with Feature and Class Frequency Statistics.
Dataset ID Dataset Name Number of Instances Number of Features Numerical Features Categorical Features Binary Features Number of Classes Min-Max Class Freq 3 kr-vs-kp 3196 37 0 37 35 2 0.91 6 letter 20000 17 16 1 0 26 0.90 11 balance-scale 625 5 4 1 0 3 0.17 12 mfeat-factors 2000 217 216 1 0 10 1.00 14 mfeat-fourier 2000 77 76 1 0 10 1.00 15 breast-w 699 10 9 1 1 2 0.53 16 mfeat-karhunen 2000 65 64 1 0 10 1.00 18 mfeat-morphological 2000 7 6 1 0 10 1.00 22 mfeat-zernike 2000 48 47 1 0 10 1.00 23 cmc 1473 10 2 8 3 3 0.53 28 optdigits 5620 65 64 1 0 10 0.97 29 credit-approval 690 16 6 10 5 2 0.80 31 credit-g 1000 21 7 14 3 2 0.43 32 pendigits 10992 17 16 1 0 10 0.92 37 diabetes 768 9 8 1 1 2 0.54 38 sick 3772 30 7 23 21 2 0.07 44 spambase 4601 58 57 1 1 2 0.65 46 splice 3190 61 0 61 0 3 0.46 50 tic-tac-toe 958 10 0 10 1 2 0.53 54 vehicle 846 19 18 1 0 4 0.91 151 electricity 45312 9 7 2 1 2 0.74 182 satimage 6430 37 36 1 0 6 0.41 188 eucalyptus 736 20 14 6 0 5 0.49 300 isolet 7797 618 617 1 0 26 0.99 307 vowel 990 13 10 3 1 11 1.00 458 analcatdata_authorship 841 71 70 1 0 4 0.17 469 analcatdata_dmft 797 5 0 5 1 6 0.79 1049 pc4 1458 38 37 1 1 2 0.14 1050 pc3 1563 38 37 1 1 2 0.11 1053 jm1 10885 22 21 1 1 2 0.24 1063 kc2 522 22 21 1 1 2 0.26 1067 kc1 2109 22 21 1 1 2 0.18 1068 pc1 1109 22 21 1 1 2 0.07 1461 bank-marketing 45211 17 7 10 4 2 0.13 1462 banknote-authentication 1372 5 4 1 1 2 0.80 1464 blood-transfusion-service-center 748 5 4 1 1 2 0.31 1468 cnae-9 1080 857 856 1 0 9 1.00 1475 first-order-theorem-proving 6118 52 51 1 0 6 0.19 1478 har 10299 562 561 1 0 6 0.72 1480 ilpd 583 11 9 2 2 2 0.40 1485 madelon 2600 501 500 1 1 2 1.00 1486 nomao 34465 119 89 30 3 2 0.40 1487 ozone-level-8hr 2534 73 72 1 1 2 0.07 1489 phoneme 5404 6 5 1 1 2 0.42 1494 qsar-biodeg 1055 42 41 1 1 2 0.51 1497 wall-robot-navigation 5456 25 24 1 0 4 0.15 1501 semeion 1593 257 256 1 0 10 0.96 1510 wdbc 569 31 30 1 1 2 0.59 1590 adult 48842 15 6 9 2 2 0.31 4134 Bioresponse 3751 1777 1776 1 1 2 0.84 4534 PhishingWebsites 11055 31 0 31 23 2 0.80 4538 GesturePhaseSegmentationProcessed 9873 33 32 1 0 5 0.34 6332 cylinder-bands 540 40 18 22 4 2 0.73 23381 dresses-sales 500 13 1 12 1 2 0.72 23517 numerai28.6 96320 22 21 1 1 2 0.98 40499 texture 5500 41 40 1 0 11 1.00 40668 connect-4 67557 43 0 43 0 3 0.15 40670 dna 3186 181 0 181 180 3 0.46 40701 churn 5000 21 16 5 3 2 0.16 40966 MiceProtein 1080 82 77 5 3 8 0.70 40975 car 1728 7 0 7 0 4 0.05 40978 Internet-Advertisements 3279 1559 3 1556 1556 2 0.16 40979 mfeat-pixel 2000 241 240 1 0 10 1.00 40982 steel-plates-fault 1941 28 27 1 0 7 0.08 40983 wilt 4839 6 5 1 1 2 0.06 40984 segment 2310 20 19 1 0 7 1.00 40994 climate-model-simulation-crashes 540 21 20 1 1 2 0.09 41027 jungle_chess_2pcs_raw_endgame_complete 44819 7 6 1 0 3 0.19
We evaluate all methods on the OpenML CC-18 benchmark. Four very large image datasets are excluded a priori due to memory issues affecting most baselines: Devnagari-Script (OpenML ID 40923), MNIST (554), Fashion-MNIST (40996), and CIFAR-10 (40927). This leaves 68 datasets as the default pool. Some methods are run on fewer datasets due to memory limits, method-specific constraints, or pretraining overlap. Table 26 reports the coverage per method and the reason for any missing datasets.
Table 26: Dataset coverage on OpenML CC-18 after excluding four image datasets (IDs: 40923, 554, 40996, 40927). Default pool = 68 datasets.
Method Used Missing Primary reason for missing datasets AutoGluon 68 0 CatBoost 68 0 LightGBM 68 0 XGBoost 68 0 RealMLP 68 0 TabM 68 0 ResNet 68 0 FT-Transformer 68 0 MLP 67 1 Memory constraint on one dataset SAINT 61 7 Memory constraints TabNet 62 6 Memory constraints CARTE 62 6 Memory constraints TabICL 68 0 TP-BERTa 15 53 Memory constraints TabPFN 17 51 Method limits: $\leq$1000 samples, $\leq$100 features, $\leq$10 classes TabPFNv2 49 19 Method limits: $\leq$10,000 samples, $\leq$500 features, $\leq$10 classes XTab 55 13 Excluded due to pretraining overlap
Appendix F Comparison with AutoML methods
In our study, we include AutoGluon Erickson et al. (2020), a prominent AutoML library, to compare against the Deep Learning methods. We consider two versions of AutoGluon: one where we perform hyperparameter optimization (HPO) and the officially recommended version configured with presets=best_quality. We compare all methods within the Deep Learning family to these versions of AutoGluon.
Figure 27: Distribution of ranks for the Deep Learning Models (13 methods) and AutoML (1 method) classifier families. Left: AutoGluon with hyperparameter optimization (HPO). Right: AutoGluon in its recommended configuration. The boxplot illustrates the rank spread, with medians represented by red lines and whiskers showing the range.
Figure 27 presents boxplots of the rank distributions for all Deep Learning methods compared to AutoGluon. The left-hand side shows results against the HPO-tuned version of AutoGluon. TabPFNv2 and TabICL achieve the best overall performance, sharing the same mean rank, with TabICL having a slightly better median rank. Both outperform all other methods, including AutoGluon (HPO). TabM also performs competitively with a median rank around 3, while most other Deep Learning methods rank lower.
The right-hand side compares against the recommended version of AutoGluon and reveals a similar picture. TabPFNv2 matches AutoGluon in both mean and median rank, though AutoGluon’s interquartile range extends slightly lower, indicating marginally better performance. TabICL attains the same mean rank as TabPFNv2 and AutoGluon, but with a better median rank.
Figure 28: Comparative analysis of Deep learning models against AutoGluon. Left: AutoGluon with hyperparameter optimization (HPO). Right: AutoGluon in its recommended configuration.
Consistent with our previous analyses, we also present critical difference (CD) diagrams to summarize the ranking comparisons. Figure 28 shows, on the left, the CD diagram comparing AutoGluon with HPO against the Deep Learning methods across all datasets, and on the right, the diagram for AutoGluon with its recommended configuration.
On the left, TabICL and TabPFNv2 achieve the best average rank of 3, followed by TabM and then AutoGluon (HPO). The diagram indicates that differences among TabICL, TabPFNv2, and TabM are not statistically significant; however, TabICL and TabPFNv2 significantly outperform all remaining methods.
On the right, TabICL is the top-ranked method, followed closely by TabPFNv2 and AutoGluon. The absence of connecting bars for these three methods indicates that their performance is statistically significantly better than that of all other Deep Learning baselines.
Figure 29: Comparative analysis of Deep learning models against AutoGluon in the small data domain (number of instances $\leq 1000$). Left: AutoGluon with hyperparameter optimization (HPO). Right: AutoGluon in its recommended configuration.
We further perform the same analysis in the small-data setting, defined as datasets with $\leq 1000$ instances, which allows us to include TabPFN in the comparison. The results are shown in Figure 29.
In the left diagram, TabPFNv2 and TabICL lead the rankings, closely followed by TabM, all outperforming the HPO-tuned version of AutoGluon. In contrast, the right diagram shows that AutoGluon with its recommended configuration ranks highest, surpassing even TabICL and TabPFNv2, although the differences are not statistically significant.
Appendix G Analysis of Dataset Characteristics: Instances and Features
To analyze the relationship between dataset size and the performance of different methods, we categorize datasets based on two key attributes: the number of instances and the number of features.
  • Instance-based Categorization:
    • Datasets with 1000 or fewer instances.
    • Datasets with 1001 to 5000 instances.
    • Datasets with 5001 to 10000 instances.
    • Datasets with 10001 to 50000 instances.
    • Datasets with more than 50000 instances.
  • Feature-based Categorization: Within each instance-based group, datasets are further divided based on the number of features:
    • Datasets with 100 or fewer features.
    • Datasets with 101 to 500 features.
    • Datasets with 501 to 1000 features.
    • Datasets with more than 1000 features.
  • Unavailable Results: Having split the datasets into these groups, we note the ones in which no dataset belongs:
    • Datasets with instances between 5001 and 10000, and features between 100 and 500.
    • Datasets with instances between 5001 and 10000, and features greater than 1000.
    • Datasets with instances between 10001 and 50000, and features greater than 1000.
    • For datasets with more than 50000 instances, we only have results for datasets with 100 or fewer features.
    • For datasets with fewer than 1000 instances, we only have results for datasets with 100 or fewer features.
For the analysis, we present boxplots and critical difference diagrams, if the number of datasets is at least 10 for meaningful analysis. If the number of datasets in a group is fewer than 10, we use tabular results instead of boxplots or critical difference diagrams.
G.1 Datasets with fewer than 1000 instances
In this section, we focus on datasets with fewer than 100 features and fewer than 1000 instances, resulting in a total of 18 datasets used in our study. Consequently, most methods in Figure 30 are evaluated on 18 datasets. However, there are a few exceptions: TabPFN and TabPFNv2 are incompatible with one dataset, ”vowel,” due to it containing more than 10 classes; XTab excludes 2 datasets that were part of its pretraining phase; and TP-Berta encounters memory limitations on 10 out of the 18 datasets, reducing its coverage.
Figure 30: Distribution of ranks for all the methods in the small data domain. The boxplot illustrates the rank spread, with medians represented by black lines, means represented by diamonds and whiskers showing the range.
Figure 30 reveals that AutoGluon achieves the strongest overall performance, closely followed by TabPFNv2 and TabICL. Among feedforward networks, TabM, MLP and RealMLP rank well, though TabM reaches a lower rank, while RealMLP has a better mean and median rank compared to MLP. Among the other dataset-specific neural networks, FT-Transformer and SAINT perform comparably. Interestingly, MLP-like methods also show a lower median rank than the classical CatBoost, LightGBM and XGBoost, although CatBoost occasionally achieves ranks as low as 2.5. By contrast, TabNet and the fine-tuning–based models generally exhibit the weakest performance.
Figure 31: Left: Distribution of ranks for all the methods in the small data domain. The boxplot illustrates the rank spread, with medians represented by black lines, means represented by diamonds and whiskers showing the range. Right: Comparative analysis of all the methods.
Similarly, Figure 31 shows a boxplot on the left, evaluated on the same datasets but excluding TP-BERTa—and a critical difference diagram on the right. A clear pattern emerges: in the small-data domain, in-context learning methods, are highly competitive, followed by dataset-specific neural architectures (e.g., TabM, MLP with PLR embeddings and RealMLP), surpassing even CatBoost, LightGBM and XGBoost.
G.2 Datasets with 1,000 to 5,000 instances
Following the previous analysis, we now examine datasets with 1,000–5,000 instances and fewer than 100 features. The results are shown in Figure 32. Similar to the small-data setting, TabICL and TabPFNv2 dominate, outperforming even AutoGluon. A notable shift, however, is the strong performance of CatBoost, which rises to fourth place overall, just behind TabM in third, while XGBoost maintains performance levels comparable to those in the smaller datasets. Furthermore, dataset-specific neural networks continue to outperform fine-tuned networks, with TabM and MLP with PLR embeddings standing out for their strong performance. Both achieve better median ranks and narrower interquartile ranges than XGBoost and LightGBM.
Figure 32: Distribution of ranks for all the methods in the datasets with 1000 to 5000 instances, and less than 100 features. The boxplot illustrates the rank spread, with medians represented by black lines, means represented by diamonds and whiskers showing the range.
Figure 33: Left: Distribution of ranks for all the methods in the common datasets with instances between 1000 and 5000, and features fewer than 100. The boxplot illustrates the rank spread, with medians represented by red lines and whiskers showing the range. Right: Comparative analysis of all the methods.
In Figure 33, we exclude TP-BERTa again to ensure a reasonable number of common datasets, resulting in a total of 19 datasets. The left plot tells a similar story, with XGBoost now achieving slightly better median rank as the MLP with PLR embeddings. The right plot presents a critical difference diagram, showing TabICL, TabPFNv2, and AutoGluon as the top-performing methods.
For the remaining dataset categorization groups, we present only tabular results due to the limited number of datasets in these categories. Table 27 provides the results for datasets with 1000 to 5000 instances and 100 to 500 features. Similarly, Table 28 summarizes the performance for datasets in the 500 to 1000 features range, while Table 29 presents results for datasets with more than 1000 features. Detailed results for all other dataset categorization groups can be found below.
Table 27: Classifier Performance for Instance Range: 1000-5000 and Feature Range: 100-500
Dataset dna mfeat-factors mfeat-pixel semeion
AutoGluon 0.995385 0.999350 0.999403 0.998506
CARTE 0.986120 0.996064 0.996175 0.993378
CatBoost 0.995028 0.998910 0.999422 0.998687
FT-Transformer 0.990937 0.999015 0.997451 0.995548
LightGBM 0.994942 0.999125 0.999131 0.997646
MLP 0.992220 0.998875 0.998674 0.997350
RealMLP 0.994111 0.999625 0.999492 0.998976
ResNet 0.992543 0.999472 0.998690 0.997689
SAINT 0.992473 0.999385 0.999217 0.997630
TabICL 0.994123 0.999808 0.999664 0.999033
TabM 0.994505 0.999700 0.999478 0.998425
TabNet 0.991448 0.998125 0.998200 0.994019
TabPFNv2 0.995658 0.999650 0.999503 0.998288
XGBoost 0.995278 0.999004 0.999378 0.998272
XTab 0.992479 0.998443 0.998642 0.997064
Table 28: Classifier Performance for Instance Range: 1000-5000 and Feature Range: 500-1000
Dataset cnae-9 madelon
AutoGluon 0.998524 0.932817
CARTE 0.990151 0.836760
CatBoost 0.996316 0.937562
FT-Transformer 0.994497 0.747391
LightGBM 0.984404 0.924095
MLP 0.996716 0.883991
RealMLP 0.997569 0.930302
ResNet 0.997106 0.605018
TabICL 0.997840 0.711538
TabM 0.998100 0.809941
TabNet - 0.630669
XGBoost 0.997454 0.932249
XTab - 0.845746
Table 29: Classifier Performance for Instance Range: 1000-5000 and Feature Range: >> 1000
Dataset Bioresponse Internet-Advertisements
AutoGluon 0.888693 0.985963
CatBoost 0.885502 0.979120
FT-Transformer 0.820159 0.974513
LightGBM 0.886734 0.980094
MLP 0.825631 -
RealMLP 0.859065 0.973810
ResNet 0.850801 0.974187
TabICL 0.885075 0.989308
TabM 0.876671 0.985640
XGBoost 0.888615 0.982276
G.3 Datasets with 5,000 to 10,000 instances
Table 30: Classifier Performance for Instance Range: 5000-10000 and Feature Range: $\leq$ 100
Dataset GPhaseSeg first-ord-TP optdigits phoneme satimage texture wall-rob-nav AutoGluon 0.936667 0.835425 0.999925 0.973342 0.993557 0.999998 0.999993 CARTE 0.798024 0.764092 0.999112 0.948702 0.988038 0.999541 0.999505 CatBoost 0.916674 0.831775 0.999844 0.968024 0.991978 0.999948 0.999990 FT-Transformer 0.895166 0.796707 0.999616 0.965071 0.993516 0.999983 0.999900 LightGBM 0.920034 0.833949 0.999818 0.965147 0.991309 0.999890 0.999968 MLP 0.911434 0.798812 0.999794 0.967617 0.992308 0.999991 0.999689 RealMLP 0.901441 0.795637 0.999968 0.966456 0.993034 0.999999 0.998720 ResNet 0.914196 0.784636 0.999927 0.963591 0.991995 0.999999 0.999042 SAINT 0.919006 0.802392 0.999841 0.960382 0.992630 0.999976 0.999844 TabPFNv2 0.936548 0.825502 0.999897 0.973546 0.995122 0.999963 0.999936 TabM 0.933828 0.818255 0.999939 0.971200 0.994291 0.999997 0.999912 TabNet 0.850596 0.774094 0.998871 0.956279 0.987482 0.999763 0.997585 TabICL 0.951408 0.834629 0.999989 0.977493 0.993373 1.000000 0.999610 XGBoost 0.916761 0.834883 0.999855 0.967421 0.992114 0.999940 0.999981 XTab 0.886960 0.798803 0.999712 0.961749 0.992918 0.999962 0.999846
Table 31: Classifier Performance for Instance Range: 5000-10000 and Feature Range: 500-1000
Dataset isolet
AutoGluon 0.999744
CatBoost 0.999389
FT-Transformer 0.998817
LightGBM 0.999401
MLP 0.998295
RealMLP 0.999635
ResNet 0.999401
SAINT -
TabICL 0.999570
TabM 0.999750
TabNet 0.998813
XGBoost 0.999488
G.4 Datasets with 10,000 to 50,000 instances
Table 32: Classifier Performance for Instance Range: 10000-50000 and Feature Range: $\leq$ 100
Dataset Phishing Adult BankMkt Elec JM1 JngChess Letter PenDigits AutoGluon 0.997572 0.931792 0.941273 0.987260 0.770272 0.999278 0.999934 0.999725 CARTE 0.994582 0.902677 0.924664 0.909407 0.728512 0.973383 0.999440 0.999468 CatBoost 0.996482 0.930747 0.938831 0.980993 0.756611 0.976349 0.999854 0.999752 FT-Transformer 0.996760 0.914869 0.938198 0.963076 0.709321 0.999975 0.999919 0.999703 LightGBM 0.997542 0.931261 0.938470 0.989807 0.753206 0.977605 0.999825 0.999735 MLP 0.996991 0.928689 0.937054 0.969201 0.715620 0.999965 0.999894 0.999705 RealMLP 0.997208 0.923327 0.937031 0.961467 0.713988 0.999774 0.999914 0.999659 ResNet 0.996975 0.913790 0.935740 0.960658 0.720444 0.999956 0.999926 0.999638 SAINT 0.996746 0.920246 0.936560 0.967012 0.719464 0.999926 0.999853 0.999782 TabICL 0.998342 0.914430 0.940210 0.970809 0.784425 0.975471 0.999957 0.999852 TabM 0.997636 0.919662 0.941872 0.968760 0.751557 0.999985 0.999943 0.999739 TabNet 0.996196 0.882450 0.887319 0.938656 0.674043 0.991981 0.999606 0.999753 XGBoost 0.997425 0.930482 0.938384 0.987790 0.759652 0.974087 0.999819 0.999703 XTab 0.996896 - - 0.966899 0.727984 0.999950 0.999859 0.999751
Table 33: Classifier Performance for Instance Range: 10000-50000 and Feature Range: 100-500
Dataset nomao
AutoGluon 0.996892
CatBoost 0.996439
FT-Transformer 0.990908
LightGBM 0.996835
MLP 0.986577
RealMLP 0.989803
ResNet 0.993048
TabICL 0.996055
TabM 0.994828
XGBoost 0.996676
XTab 0.992727
Table 34: Classifier Performance for Instance Range: 10000-50000 and Feature Range: 500-1000
Dataset har
AutoGluon 0.999958
CatBoost 0.999941
FT-Transformer 0.999685
LightGBM 0.999959
MLP 0.999783
RealMLP 0.999959
ResNet 0.999921
TabICL 0.999913
TabM 0.999966
TabNet 0.999515
XGBoost 0.999960
G.5 Datasets with more than 50,000 instances
Table 35: Classifier Performance for Instance Range: >> 50000 and Feature Range: $\leq$ 100
Dataset connect-4 numerai28.6
AutoGluon 0.934636 0.530150
CARTE - 0.514361
CatBoost 0.921050 0.529404
FT-Transformer 0.901170 0.530315
LightGBM 0.932440 0.529077
MLP 0.927373 0.525920
RealMLP 0.928258 0.529534
ResNet 0.933333 0.528012
SAINT - 0.525822
TabICL 0.897904 0.526838
TabM 0.941654 0.529336
XGBoost 0.931952 0.529457
XTab - 0.528062
Appendix H Cost vs. Efficiency Relation of Various Model Families
To observe what is the cost vs. efficiency relation of various model families, we plot the intra-search space normalized Average Distance to the Maximum (ADTM) Wistuba et al. (2016) in Figure 34, illustrating how quickly each method converges to its best solution during the HPO process.
The plot shows that XGBoost is the fastest, reaching nearly optimal performance within just 5 hours. The ResNet and MLP architecture also demonstrate notable speed, followed closely by CatBoost.
Figure 34: Intra-search space normalized average distance to the maximum over cumulative training time (seconds).
Figure 35: Left: Performance profiles based on inference time. Right: Performance profiles based on total time. Steeper curves indicate better overall performance and efficiency across datasets.
Overall, the gradient boosting methods (GBDTs) converge faster than the deep learning models. XTab, which shares the same transformer architecture as FT-Transformer, exhibits quicker convergence, likely due to its static architecture, while the FT-Transformer’s architectural components were also tuned. On the other hand, TP-BERTa is the slowest to converge, likely due to the high computational demands of its BERT-like architecture.
In Figure 35, we show the performance profiles of the models considered. We first normalize the performance values and the logarithmic time values.
Formally, let $P_{i}^{(j)}$ denote the performance of an algorithm $i$ on dataset $j$, and $T_{i}^{(j)}$ the corresponding executing time. We define $m_{\ast}^{(j)}=\max_{i}~P_{i}^{(j)}$ to be the best performance achieved across all algorithms, and $t_{\dagger}^{(j)}=\max_{i}~T_{i}^{(j)}$ as the longest runtime observed. Next, we compute for each algorithm $i$ and dataset $j$ the performance gap $\operatorname{gap}_{i}^{(j)}=(m_{\ast}^{(j)}-P_{i}(j))/m_{\ast}^{(j)}$ and the temporal gain $\operatorname{tgain}_{i}^{(j)}=(t_{\dagger}^{(j)}-T_{i}^{(j)})/t_{\dagger}^{(j)}$. Using these, we define the Performance-Time Ratio $\operatorname{ptr}_{i}^{(j)}=\operatorname{gap}_{i}^{(j)}/\operatorname{tgain}_{i}^{(j)}$ quantifying the trade-off between performance loss and time savings. To further enable comparison across datasets, we normalize the PTR values to the range $[0,1]$, such that values closer to 1 (0) indicate a better (poorer) performance-time trade-off. When computing the cumulative distribution of these normalized values, we count how many PTR values are less than or equal to a threshold $\tau\in[0,1]$ for each algorithm $i$. Hence, the resulting plot indicates how often an algorithm achieves favorable performance-time trade-offs. Curves that rise more steeply and reach higher proportions for lower $\tau$ values correspond to better overall performance-time characteristics. A red dashed corner frame in the top-left highlights the optimal trade-off.
On the left, the performance profiles are shown w.r.t. the measured inference time. The evaluation shows that feed-forward neural network models yield high performance-time ratios. TabM leads the performance-time ratios, followed by AutoGluon, RealMLP, and CatBoost. Although the AutoML framework AutoGluon shows strong performance values as discussed in more detail in Appendix F, this entails a high computational burden resulting in increased temporal costs. Both TP-Berta and TabPFN are only evaluated on a small subset of the available datasets as reflected in their low cumulative dataset coverage, where the latter shows its strong performance on the datasets it has been evaluated on by a steep increase for low values of $\tau$. The approaches SAINT, TabNet, CARTE, and XTab lie in an intermediate range, where CARTE and SAINT show slightly better performance-cost ratios compared to AutoGluon with increased values for the performance ratio $\tau$, i.e., considering a broader range of various datasets. The right plot shows the equivalent performance profiles w.r.t. the measured total time. Notably, XGBoost, AutoGluon with hyperparameter optimization, and TabM achieve strong performance-time ratios. Transformer-based models are outperformed by more lightweight models like CatBoost and ResNet which both show competitive results. From the model family encompassing foundation models, XTab shows the strongest performance. However, it is outperformed by classical GBDT approaches. TabPFN, an in-context learning model, is only applicable on small data regimes, and is therefore not competitive across the full benchmark suite.
Figure 36: Performance profiles in the small data domain. Left: Performance profiles based on inference time. Right: Performance profiles based on total time. Steeper curves indicate better overall performance and efficiency across datasets.
Figure 37: Performance profiles in the large data domain. Left: Performance profiles based on inference time. Right: Performance profiles based on total time. Steeper curves indicate better overall performance and efficiency across datasets.
Small-Data Domain.  In Figure 36, the performance profiles are shown w.r.t. the measured inference time (left) and the measured total time (right) in the small-data regime. The models CatBoost and AutoGluon yield the best performance-time ratios, with SAINT as a transformer-based model and TabM being competitors when increasing the performance ratio $\tau$. The models FT-Transformer, ResNet, and TabNet yield similar results, where amongst these models FT-Transformer performs slightly better for small performance ratios. All of them yield a better performance-cost ratio for a larger amount of datasets. TB-BERTa exhibits the least favorable trade-off as the inference time largely outweighs the performance.
On the right side, the performance plots are given w.r.t. the total time. In the small-data regime, the model TabPFN yields strong performance-cost ratios resulting in a superior performance followed by XGBoost and the feedforward models MLP and ResNet. Due to the larger training time, the fine-tuned models CARTE, TabNet, and TP-BERTa do not match the performance ratios of models from other architectural classes.
Large-Data Domain.  In Figure 37, the performance profiles are shown w.r.t. the measured inference time (left) and the measured total time (right) in the large-data regime. As discussed in Section 5, TabPFN is only applicable to small-data, hence, it is not included in the large-data analysis. Regarding the inference time, TabM is superior to all other competitors, followed by the models AutoGluon as an AutoML-driven approach and CatBoost from the GDBTs. FT-Transformer shows strong results on about half the datasets used in our analysis, but fails to maintain performance across the entire benchmark suite and is comparable to the models TabNet, CARTE, and SAINT. This group of models shows slightly better trade-off values compared to other competitors for an increase performance ratio $\tau$. Like before, TP-Berta struggles to be competitive and shows the worst performance-cost ratios.
When considering the total amount of time, the models AutoGluon (HPO) and XGBoost show the strongest performance-cost trade-offs. It is followed by CatBoost from the GDBTs family, and the lightweight feedforward networks, ResNet and MLP. From the fine-tuned models, XTab beats CARTE, whereas FT-Transformer wins over SAINT and TabNet from the transformer-based approaches. TP-BERTa is not competitive with any of the other approaches.
Appendix I Influence of meta-feature characteristics on the predictive performance
Following the methodology of McElfresh et al. (2023), we employed the PyMFE library (Alcobaça et al., 2020) to extract meta-features from the datasets used in our study. Specifically, we extracted General, Statistical, and Information-theoretical meta-features.
Figure 38 displays the mean correlation coefficients of the most significant meta-features concerning the performance of all methods, averaged across datasets. To produce this plot, we first calculate the correlation coefficients between each method’s performance and each meta-feature for all datasets. For each method, we then selected the top $k$ meta-features with the highest absolute value of the correlation coefficients across all datasets, identifying them as the most important ones for that specific method. We compiled a list of significant meta-features by taking the union of these top meta-features across all methods.
Figure 38: Mean correlation coefficient of most important meta-features with performance across all methods
For each meta-feature in this combined list, we computed the mean of its correlation coefficients across datasets for all methods. Figure 38 illustrates that TabPFN and TPBERTa significantly deviate from the overall pattern observed in the other methods, exhibiting negative correlations for the meta-features mad.mean, median.sd, t_mean.mean, and t_mean.sd. To determine whether this deviation is due to the inherent properties of these methods or is a consequence of the limited number of datasets they were evaluated on, we repeated the analysis for all methods using only the datasets on which TabPFN and TPBERTa were run.
Figure 39: Mean correlation coefficient of most important meta-features with performance across all methods on datasets with results for TabPFN and TPBERTa
Figure 39 demonstrates that when the analysis is confined to only the intersection of datasets on which TabPFN and TPBERTa were evaluated, the previously observed deviation disappears. This suggests that the initial divergence was likely due to the limited number of datasets rather than the inherent properties of these methods. Therefore, it appears that all methods, regardless of their method families, are similarly influenced by the meta-features in terms of their predictive performance. In general, the strongest correlation coefficients are observed for three meta-features: eq_num_attr, w_lambda, and can_cor.mean.
The eq_num_attr meta-feature, which measures the number of attributes equivalent in information content for the predictive task, exhibits a strong negative correlation with performance across most methods. This suggests that methods generally perform worse on datasets with high feature redundancy, likely due to challenges in handling overlapping information or overfitting. Similarly, the w_lambda meta-feature, which computes Wilk’s Lambda to quantify the separability of classes in the feature space, also shows a consistently negative correlation. This indicates that methods struggle on datasets with poor class separability, where the features do not adequately distinguish between the target classes. Conversely, the can_cor.mean meta-feature, representing the mean canonical correlation between features and the target, shows a positive correlation with performance. This implies that methods perform better on datasets where the features are strongly predictive of the target variable, highlighting their reliance on well-aligned feature-target relationships.
Generally, the findings align with the common intuition of the performance of ML methods under sub-optimal class separation and further validate the empirical protocol of our study. For detailed explanations of the meta-feature abbreviations used in the plots, please refer to the official PyMFE documentation [^14] https://pymfe.readthedocs.io/en/latest/auto_pages/meta_features_description.html.